### 21 Jan 2024 v3

# FedDRL: A Trustworthy Federated Learning Model Fusion Method Based on Staged Reinforcement Learning

Abstract
Traditional federated learning uses the number of samples to calculate the weights of each client model and uses this fixed weight value to fusion the global model. However, in practical scenarios, each client’s device and data heterogeneity leads to differences in the quality of each client’s model. Thus the contribution to the global model is not wholly determined by the sample size. In addition, if clients intentionally upload low-quality or malicious models, using these models for aggregation will lead to a severe decrease in global model accuracy. Traditional federated learning algorithms do not address these issues. To solve this probelm, we propose FedDRL, a model fusion approach using reinforcement learning based on a two staged approach. In the first stage, Our method could filter out malicious models and selects trusted client models to participate in the model fusion. In the second stage, the FedDRL algorithm adaptively adjusts the weights of the trusted client models and aggregates the optimal global model. We also define five model fusion scenarios and compare our method with two baseline algorithms in those scenarios. The experimental results show that our algorithm has higher reliability than other algorithms while maintaining accuracy.
1 Introduction
With the development of deep learning technologies, numerous industries are applying them to their respective fields, resulting in the emergence of intelligent transportation, smart logistics, and smart healthcare models. These technologies reduce the industry’s production and management costs, increase operational efficiency, and accelerate the industry’s digitization process. However, most current deep learning-based model training uses a supervised learning method. The number and diversity of samples is the key to training a high-quality model. Therefore, how to obtain massive and diverse data samples has become the first step to training deep learning models. This method leads to the source of these samples being no longer limited to a single industry but to multiple industries collaborating to construct large-scale datasets. Moreover, with the emergence of cross-industry collaborative data analysis models, data analysis sources are no longer limited to a single organization but involve multiple parties. Although the participation of multiple parties in the data analysis process can solve the problem of cross-domain data analysis, this strategy also faces a new challenge. The traditional data analysis method requires uploading each participant’s raw data to a central node and training models on the central node. However, this workflow involving raw data transfer poses security and privacy risks. Meanwhile, raw data transmission is not permitted for data-sensitive industries, such as hospital patient and financial industry data. How to achieve federated data analysis of multiple parties while protecting data security is becoming hot research.
Federated learning is an emerging paradigm for distributed computing. The structure of federated learning has a central node and multiple participants, where multiple participants train models with their private data, and the central node achieves knowledge extraction and information fusion of each participant by fusing the models uploaded by each client. Since the process does not directly use the participants’ data, federated learning can achieve joint data analysis of multiple participants while protecting user privacy in advance. To solve mobile device word input privacy protection issues, Google proposed federated learning and the FedAvg method [1], which ensure user privacy by transmitting models instead of raw data. Increasingly, companies are applying federated learning in various domains to tackle cross-industry data analysis challenges. However, in practical federated learning environments, the heterogeneity of data among participants gives rise to the non-iid problem, which makes model convergence challenging during the fusion stage of the FedAvg method. Additionally, variations in computational resources among participants lead to slow convergence of the global model. To address these challenges, some research has optimized the global model convergence problem by adding control parameters in the client-side model fusion process, such as FedProx [2], Scaffold [3], and FedNova [4]. The Moon algorithm [5] optimizes the convergence of the global model by introducing a comparative loss between the client-side model and the global model. There is also some research [6] [7] to improve the accuracy of the global model by selecting the clients involved in the global model. These algorithms somewhat solve the global model’s convergence problem in the non-iid data among participants.
Current traditional federal learning methods, such as FedAvg and FedProx, use a ratio based on the number of samples of each participant to the total number of samples to calculate the model weights of each participant. However, in a practical scenario, the method has two problems. On the one hand: the method does not consider the different participants’ model quality, which gives higher weights to the client with a larger number of samples. However, due to the variability of computational resources and sample quality across clients, it does not mean that a client can train a higher-quality model when a client has more samples. Therefore, using these algorithms assumes that the sample quality and computational resources are consistent across clients. On the other hand, if the participants have malicious attack behavior and intentionally upload malicious models to participate in fusion, which will seriously influence the global model convergence process. Therefore it becomes increasingly important to design adaptive defenses for federated learning systems [8]. How to eliminate malicious models in federated learning model fusion has become a vital issue. At the same time, when the client uploads low-quality models to participate in fusion, how to adaptively adjust each model’s fusion weights based on the models’ quality is also an urgent problem to be solved. Some research has used reinforcement learning methods to address these issues of weighting. Such as, Favor [7] method uses the DDPG to calculate the weights of the participant models. Some research has also used reinforcement learning to solve the problems of device selection [9] [10], resource optimization [11] [12], and communication optimization in IoT federated learning scenarios.
Reinforcement learning (RL) methods use a try-and-learn strategy. The core of the reinforcement learning approach is to train an intelligent agent. The agent uses different actions to interact with the external environment. The environment provides reward and punishment values based on the selected actions of the intelligence. The intelligent body achieves optimal action selection by maximizing the reward value. However, when we use reinforcement learning, we must address some issues. On the one hand, we must collect training samples by continuously interacting with the environment. When the cost of interacting with the external environment is too high, or the interaction cost is unacceptable (e.g., driverless scenario), how to collect samples quickly will directly affect the training time of reinforcement learning. On the other hand, when the action space of the agent is high-latitude and continuous, this will lead to an increasing sampling time. Therefore, due to these two problems, traditional single-agent reinforcement learning training methods can take much time. When we apply reinforcement learning to federated learning, we must solve those problems. As the number of participants grows, the time cost of training agents will increase. How to optimize the action space for reinforcement learning and speed up the training process of the agent is an urgent problem to be solved.
To address these issues, we propose a staged reinforcement learning algorithm (FedDRL) to solve the problem of model weight calculation in federated learning. The algorithm consists of two phases: the selection of trusted clients and the adaptive assignment of weights. We use a staged approach to accelerate the agent’s training. We use a distributed reinforcement learning algorithm to speed up the collection of training samples. The contributions of this paper is as follows.
  • We propose a federated learning model fusion method based on reinforcement learning, which solves the problem of malicious clients’ participation in model fusion and clients’ model weight assignment.
  • We propose a trusted client selection strategy that eliminates malicious models from participating in global model fusion while enabling the selection of trusted clients.
  • We propose a adaptive weight assignment method that adaptively adjusts the weights corresponding to the clients according to the quality of their uploaded models and therefore computes the optimal global model.
  • We present five types of model fusion scenarios to validate the performance of each algorithm. We also compare the performance of our algorithm with the baseline algorithm on three public datasets.
2 Related Work
2.1 Federated learning
Research of federated learning focuses on solving three problems: how to improve generalization of the global model on the server side, how to personalize the model on the client side.Therefore, Federated learning can be divided into server-side and client-side optimization algorithms. Google first proposed the FedAvg method [1] to solve the server-side global model fusion problem . However, the client in the actual scenario has device and sample heterogeneity problems, which can lead to global model convergence problems in the FedAvg method. To optimize the global model convergence problem, Karimireddy et al. proposed the Scaffold algorithm [3], which accelerates the convergence of the global model by introducing a control variable to correct the client-side drift problem. Li et al. proposed the FedProx algorithm [2], which implements the correction of deviated clients by adding the regularization function to the client model. Wang et al. proposed the FedNova algorithm [4], which solves the global model convergence problem by normalizing the parameters on both the client and server sides. Li et al. also proposed the MOON [5] method based on model comparison learning, which solves the global model convergence problem by calculating the similarity between the global model and the client model and continuously decreasing the distance representation of both to achieve the training of the client-side model.
2.2 The challenges of Non-IID Data Distribution
The Non-IID problem is the key that leads to the convergence of the federated learning model. Zhao et al. study the performance of different federated learning methods on the Non-IID dataset and experimentally demonstrate that the Non-IID problem poses a significant challenge to the accuracy of the federated learning algorithm [13]. Therefore some research focuses on solving the Non-IID problem of federated learning. For example, Zhang et al. proposed the FedPD method [14], which offers strategies that can handle non-convex objective functions to optimize models and communication. In addition, Gong et al. proposed the AutoCFL method [15], which adaptively adjusts the number of training rounds for the client while introducing weighted voting of the client clustering strategy to eliminate the effects of Non-IID and imbalanced data. Huang et al. proposed the FedAMP method [16], which solves the client-side model personalization problem caused by Non-IID data by implementing a personalized update of the client-side model. Li et al. proposed the Fedbn method [17], which adds a batch normalization layer to the local model to solve the problem of feature shift in data heterogeneity. Briggs et al. proposed a hierarchical clustering approach (FL+HC)[18], which optimizes the model’s performance on Non-IID datasets by training clients in groups and then having each group trained independently on a dedicated model. Furthermore, Gao et al. proposed the Feddc method [19], which addresses the data heterogeneity problem by bridging the gap between client and global model parameters through a control variable. Finally, mu et al. proposed the Fedproc method [20], which guides the client model training by introducing a loss of comparison between the client model and the global model, thus enabling the global model to perform well on Non-IID data to achieve good performance.
2.3 Federated reinforcement learning
Since reinforcement learning has the advantage of adaptive learning, some research applies reinforcement learning to federated learning. Some research focuses on improving the performance of global models based on reinforcement learning methods. Such as Wang et al. proposed the Favor method [7], which adaptively selects clients participating in the model fusion. Sun et al. also proposed the PG-FFL framework [6], which can solve the problem of calculating client weights during model fusion. Some work uses reinforcement learning to solve device optimization problems in a federated IoT environment. such as Zhang et al. uses the DDPG method [9] to achieve the selection of optimal devices. Zhang proposed the FedMarl algorithm [21], which uses Multi-agent reinforcement learning to achieve node selection in federated learning. Yang et al. also proposed a digital twin architecture (DTEI) [10], which uses reinforcement learning methods to solve the device selection problem. Some research focus on solving the resource optimization and scheduling problem in IoT environment. Zhang et al. proposed the RoF approach[11], which achieves optimal resource scheduling based on multi-intelligent reinforcement learning. Rjoub et al. designed a trusted device selection method [22] based on a reinforcement learning approach to achieve resource scheduling. Rjoub et al. proposed the DDQN-Trust method [12], which uses the Q-learning approach to evaluate the credit scores of devices to achieve optimal scheduling. To optimize the communication problem of federated learning, Yang et al. proposed a reinforcement learning-based model evaluation method [23] to participate in model training and model fusion by selecting the optimal devices. Most current research uses reinforcement learning to solve federated learning problems in IoT environments. These include optimal device selection, resource optimization, and communication optimization. Some work uses reinforcement learning methods to solve the client-side weight assignment problem in federated learning. However, most of these methods are based on single-agent training methods and need to address the training time overhead of reinforcement learning.
3 Method
3.1 Problem Definition
In this section, we first analyze the problems of the current federated learning approach and then propose a solution to the problem. In the federation learning process, we aim to get a higher-accuracy global model by aggregating all clients’ local models through the server side.
In the federated learning environment, we define n clients are involved in model fusion, and the client is denoted as $C_{i}$ where $C_{i}\in\left\{C_{1},C_{2},C_{3}\ldots C_{n}\right\}$. Each client has a network model $M_{i}$, where $M_{i}\in\left\{M_{1},M_{2},M_{3}\ldots M_{n}\right\}$. Each client has its private data $D_{i}$, where $D_{i}\in\left\{D_{1},D_{2},D_{3}\ldots D_{n}\right\}$. The number of samples in each dataset is $S_{i}$, where $S_{i}\in\left\{S_{1},S_{2},S_{3}\ldots S_{n}\right\}$. The total number of samples is $\sum_{i=1}^{N}S_{i}$. We define the $\theta_{i}$ as a model parameter of $M_{i}$. where $\theta_{i}\in\left\{\theta_{1},\theta_{2},\theta_{3}\ldots\theta_{n}\right\}$.
Also, we define the server-side model aggregation process for each round as equation 1.
$\theta_{\text{global }}=\sum_{i=1}^{N}w_{i}\theta_{i},\text{ where }w_{i}=\frac{S_{i}}{\sum_{i=1}^{N}S_{i}},\quad w_{i}\geq 0,\sum_{i=1}^{N}w_{i}=1$ (1)
Where $w_{i}$ is the weight of each client model parameter. The traditional method of calculating this weight in federal learning is determined by the ratio of the number of samples of the client to the total number of samples. This means that the larger the number of samples corresponding to the client, the larger the weight assigned. However, in the actual environment, due to the heterogeneity of the client data and the computing resources, it is impossible to determine the weights by the ratio of the number of samples for each client. Therefore, we need to solve two problems.
Problem 1: The sample size of a client accounts for 20% of the total samples, but the accuracy of this client model is only 53%, which will reduce the accuracy of the global model if the traditional federated model fusion algorithm is used.
Problem 2: If the client involved in the model fusion has malicious attacks and intentionally modifies the model output to make its model accuracy 10%, the global model accuracy will be seriously affected if the malicious models are selected for fusion using the traditional federated model fusion algorithm.
To solve these two problems, we need an adaptive weight calculation method that can set the weight of the malicious model to 0, which means that malicious models are prohibited from participating in the fusion process. At the same time, we realize the adaptive assignment of weights to each client and try to increase the weights of high-quality client models so that the global model gets the optimal accuracy rate.
If we use the single-agent reinforcement learning approach to solve the above problems, we face another issue again. On the one hand, the action space of the agent can be huge as the number of fused clients increases, which also leads to a long time required to train the agent model. On the other hand, a single agent method can only interact with a single environment, and that sampling process takes a long time. We solve these two problems in two steps by borrowing the idea of hierarchical reinforcement learning, which can solve the problem of the long training time of reinforcement learning. These two stages are selecting trusted clients and the optimal weight assignment process.
Stage 1: In this phase, We need select K-trusted models from N models to participate in the global model fusion. However, we do not know which client uploaded a malicious model. To solve the problem, We use reinforcement learning to achieve dynamic selection and autonomous screening of client models as equation 2.
$\left\{M_{a},M_{\mathrm{b}},\ldots M_{\mathrm{k}}\right\}\leftarrow\text{ SelectTrustedModel }\left(\left\{M_{1},M_{2},\ldots M_{\mathrm{n}}\right\}\right)$ (2)
Stage 2: Based on step 1, we assign optimal weights to the trusted models to improve the accuracy of the global model and represent the process as equation 3.
$\left\{W_{1},W_{2},\ldots W_{\mathrm{n}}\right\}\leftarrow\operatorname{AdaptCalculateWeight}\left(\left\{M_{a},M_{b},\ldots M_{\mathrm{k}}\right\}\right)$ (3)
$AdaptCalculateWeight(.)$ is a weight-adaptive computation method. $W_{i}$ is the optimal computational weight for each client of the output.
Why do we use a phased approach to reinforce learning? Let’s use an example to illustrate the reason for using it. We need a robot to learn the cooking process using a reinforcement learning approach. The cooking steps are divided into three stages: washing, chopping, and cooking. The robot needs to learn the best actions to cook a good dish. Traditional reinforcement learning is to find the best action that satisfies all three stages at the same time. However, if the robot does not wash the vegetables well, it cannot cook a good dish in the subsequent stages either. Therefore, it is a complete waste of time to try the next step of action if the optimal action of the previous stage is not learned. If the robot learns in multiple stages, it first learns the optimal action in the cleaning process and then proceeds to the next step of learning the optimal action. This will save time and make it easier to get the best solution.
Therefore, if the malicious models are not screened and the weight assignment method is used for malicious models, it will increase the cost of trial and error time in the reinforcement learning process. Therefore, we screen the trusted clients for weight assignment by a staged approach. This can exclude the trial and error cost of malicious models involved in weight assignment, thus optimizing the action space. Also, we use distributed reinforcement learning methods to increase the sampling speed and thus reduce the time overhead of training reinforcement learning. We will describe the implementation of each step in detail in the next section.
3.2 Trusted Client Selection Method
We propose the FedDRL approach to solve the two problems in the previous section. In the first stage, we select the trusted clients to achieve the fusion of the models by excluding the malicious clients. In the second stage, we compute the optimal model weights to improve the accuracy of the global model. The flow of the FedDRL method is shown in Figure 1.
We select trusted clients using the A2C algorithm. A2C is a synchronization algorithm achieved by improving the A3C algorithm [24]. The selection process of a trusted client for each communication round of federated learning can be treated as a Markov decision process. The reinforcement learning modeling in this phase is described as follows.
State: For the m-th communication round, we collect n models uploaded by each client and select k clients to participate in the fusion of global models. We define the $S^{t}$ as a select k client to fusion model in state t, which is shown as Equation 4, and the state includes the selected client models, the accuracy of each model, and the accuracy of the global model aggregated out by the selected client.
$S^{t}=\left\{\theta_{1}^{t},\theta_{2}^{t},\theta_{3}^{t}\ldots\theta_{k}^{t},acc_{1}^{t},acc_{2}^{t},acc_{3}^{t}\ldots acc_{k}^{t},acc_{global}^{t}\right\}$ (4)
Action: Our main objective is to eliminate the malicious models and select the trusted models to participate in the fusion. To achieve this purpose, we randomly choose several models to fusion the global model and evaluate the accuracy of global models. If a malicious model is selected to participate in the fusion, the accuracy of the global model will be very low. If only trusted clients are selected to participate in the fusion, the accuracy of the global model will also increase. We define the action of selecting a model as $a^{t}\in\left\{0,1\right\}$, where one means that the model is selected and 0 means that the model is not selected. Thus, n models correspond to the action space that can be expressed as equation 5.
$A^{t}=\left\{a_{1}^{t},a_{2}^{t},a_{3}^{t}\ldots a_{n}^{t}\right\},\quad a^{t}\in\left\{0,1\right\}$ (5)
Reward: We define the accuracy of selecting all models aggregated as $Acc_{all}$; we take this accuracy as the baseline and also represent the global model accuracy $Acc_{m}$ obtained by the m-th randomly selected client models, and calculate the reward value $Reward_{1}$ by the difference of $Acc_{m}-Acc_{all}$. Meanwhile, to avoid the intelligent body only, It can be realized that the agent selects all trusted clients as much as possible. The higher the number of models the agent sets, the higher the value of this reward. Therefore, the total reward is defined as equation 6.
$Reward=\begin{cases}\alpha*(Acc_{m}-Acc_{all})+\beta*(Number),Acc_{m}>Acc_{all}\\ \alpha*(Acc_{m}-Acc_{all}),Acc_{m}<Acc_{all}\end{cases}$ (6)
Where $\alpha$ and $\beta$ are fixed values that can be custom adjusted to balance the importance of Reward1 and Reward2, preventing one of the rewards from being too large and affecting agent training.
The agent’s goal is to obtain the maximum long-term reward value based on the discount factor $\gamma$, and the process is expressed as equation 7.
$R(\tau)=\sum_{\mathrm{t}=1}^{\mathrm{T}}\gamma^{t}\mathrm{r}_{\mathrm{t}}$ (7)
Since we use an A2C distributed reinforcement learning-based approach, as shown in Figure 1, the A2C architecture consists of a central node and K workers. Each worker contains an Actor and a Critic network, where the Actor-network generates action A, and the Critic network evaluates the action and gives the corresponding reward. Meanwhile, each worker independently interacts with the related environment to achieve sampling and training of the Actor and Critic networks. In addition, the Actor and Critic networks of the central node are used to synchronize the network information of each worker and to achieve the fusion and sharing of network parameters of multiple workers.
Figure 1: The Algorithm flow of FedDRL
Therefore our main objective is to train Actor and Critic networks. We define the Actor-network parameters as $\pi(\theta)$ and the Critic network parameters as $V(w)$. Our working process for each worker and the central node is as follows. The process is shown in algorithm 1.
Step 1: First, each worker initializes the local network by pulling the global network model parameters. Then, both networks train their respective Actor and Critic networks by interacting with the environment independently. Finally, the trained networks are uploaded to the central node.
Step 2: After the central node collects the network parameters uploaded by all workers, it updates the global model by weighted averaging. Finally, it sends it to each worker to update the network parameters.
Steps 1 and 2 are repeated according to the total number of times to obtain the final global model.
The training process of the step 1 neutralization network is as follows. The gradient of the primary communication algorithm of the policy network is calculated as equation 8.
$\nabla_{\theta}J(\theta)=\nabla_{\theta}\log\pi\left(a_{t}\mid s_{t};\theta\right)A\left(s_{t},a_{t};\mathrm{w}\right)$ (8)
Where $A\left(s_{t},a_{t};\theta_{v}\right)$ is the advantage function. The k-step sampling strategy is used in the A2C algorithm to calculate the advantage function, so the definition is expressed as equation 9.
$A\left(s_{t},a_{t};\theta_{v}\right)=\sum_{i=0}^{k-1}\gamma^{i}r_{t+i}+\gamma^{k}V\left(s_{t+k};\mathrm{w}\right)-V\left(s_{t};\mathrm{w}\right)$ (9)
The Loss function of the Actor-network is calculated as in equation 10, and The Critic network is calculated as in equation 11.
$\nabla_{\theta}J(\theta)=\nabla_{\theta}\log\pi\left(a_{t}\mid s_{t};\theta\right)\left(\sum_{i=0}^{k-1}\gamma^{i}r_{t+i}+\gamma^{k}V\left(s_{t+k};\mathrm{w}\right)-V\left(s_{t};\mathrm{w}\right)\right)$ (10)
$\nabla_{w}J(w)=\nabla_{w}\left(\sum_{i=0}^{k-1}\gamma^{i}r_{t+i}+\gamma^{k}V\left(s_{t+k};\mathrm{w}\right)-V\left(s_{t};w\right)\right)^{2}$ (11)
We update the Actor and Critic network parameters using the derivative formula as equation 12.
$w\leftarrow w+\nabla_{w}J(w),\quad\theta\leftarrow\theta+\nabla_{\theta}J(\theta)$ (12)
Finally, the network parameters of the central node are calculated by the weighted average method. The process is equation 13.
$w_{global}=\frac{1}{n}{\textstyle\sum_{1}^{n}}w_{i},\theta_{global}=\frac{1}{n}{\textstyle\sum_{1}^{n}}\theta_{i},i\in[1,n]$ (13)
Algorithm 1 FedDRL Stage 1 Algorithm Process
1: Client Models $\left\{m_{1}^{t},m_{2}^{t},m_{3}^{t},\ldots m_{\mathrm{n}}^{t}\right\}$, Round T, Worker Number K, Sampling Step Length S
2: Chosen Credible Client Model List M = $\left\{m_{2}^{t},m_{3}^{t},\ldots m_{\mathrm{k}}^{t}\right\}$
3:/* Each Worker Training Step */
4:worker $(\theta,w)\leftarrow GetGlobalParamter\left(\theta_{\text{global }},w_{\text{global }}\right)$
5:the Client Upload Current Epoch Model, Turn to State $\mathrm{s}_{0}$, $t_{\text{start }}=t=1$
6:for $e$ from 1 to $S$ do
7:     According to Current State $\mathrm{s}_{0}$ Randomly Choose Action $\mathrm{s}_{t}$
8:     ${s_{t},a_{t},r,s_{t+1}}\leftarrow\operatorname{Step}\left(\mathrm{a}_{t}\right)$ // Execute Action $\mathrm{a}_{t}$ to Acquire Reward $\mathrm{r}$ and Next State $\mathrm{s}_{t+1}$
9:     $\mathrm{t_{start}}$ = $\mathrm{t_{start}}$ + 1
10:     if $\mathrm{s_{t}}$ != terminal: R $\leftarrow$ V$(\mathrm{s_{t}};w)$ else: R = 0
11:for  $\mathrm{i}\in\left\{t-1,\ldots,\mathrm{t}_{\text{start }}\right\}$ do
12:     R $\leftarrow{\mathrm{r_{i}}+\gamma}$R // Compute Target TD
13:     $\nabla_{\theta}J(\theta)=\nabla_{\theta}\log\pi_{\theta}\left(a_{t}\mid s_{t}\right)\left(R-v\left(s_{i};w\right)\right)$ // Compute Strategy Gradient
14:     $\nabla_{w}J(w)=\nabla_{w}\left(R-v\left(s_{i};w\right)\right)^{2}$ // Compute Critic Network Gradient
15:     Update Actor Network Parameters: $\theta\leftarrow{\theta+\nabla_{\theta}J(\theta)}$
16:     Update Critic Network Parameters: $w\leftarrow w+\nabla_{w}J(w)$
17:/* Center Node Process */
18:for $round$ from 1 to $T$ do
19:     for $\mathrm{worker}_{i}$ from 1 to $K$ do
20:         Receive Each Worker Parameters$(\theta,w)$
21:         $\operatorname{Global}\left(\theta_{\text{global }},w_{\text{global }}\right)\leftarrow\mathbf{Agg}\left(\left\{\left(\theta_{1},w_{1}\right),\left(\theta_{2},w_{2}\right),\ldots\right\}\right)$ // Aggregate Parameters
22:         $\mathrm{worker}_{i}\leftarrow{SendGlobal\left(\theta_{\text{global }},w_{\text{global }}\right)}$ // Send New Parameters to Worker
23:/* Results Process */
24:Output Trusted Client Model List M = $\left\{m_{1}^{t},m_{2}^{t}\ldots m_{k}^{t}\right\}$
3.3 Adaptive weights calculation method
In this phase, we implement a distributed reinforcement learning approach based on TD3 [25]. In this phase, our main objective is to achieve the optimal weight assignment for each model. Since the weight values of the client models correspond to a continuous action space, we use the TD3 algorithm to accomplish this process. For each communication round in federation learning, we assume that K plausible client models are screened in the previous stage, where the parameters of the client models are set. The number of samples corresponding to the clients is defined as. These two terms are fixed in each federal communication round. We need to assign the weight values corresponding to each model to the k models based on the quality of the models, which is defined. We validate the accuracy Acc of each model based on the retained data. We need to train the reinforcement learning intelligences in each communication round of federation learning and achieve the global model fusion by assigning the optimal model weights by this intelligences. The process is described as follows.
(1) We define the t-th time, the action adopted by the agent to assign the weights of each model as equation 14.
$\mathrm{W^{t}}=\left\{w_{1}^{t},w_{2}^{t},w_{3}^{t}\ldots w_{k}^{t}\right\}$ (14)
(2) We aggregate the global models based on the model weights assigned by the agent, and the process is expressed as equation 15.
$\theta_{global}^{k}=\sum_{\mathrm{i}=1}^{\mathrm{t}}w_{i}^{t}\theta_{i}$ (15)
(3) We use the global model aggregated in step 2 to calculate the accuracy of this model and the reward value based on the accuracy rate. And the reward value is returned to the intelligence. The intelligence is trained by steps 1-3 until the reward value of the intelligence is stable. Finally, the optimal weights of each model are output by the intelligent body. We model the process as follows.
State: We define the state of the teeth reinforcement learning as including the number of data samples S for each client, the accuracy Acc of each client’s model, and the global model Acc based on this aggregation. The process is as in 16.
$\mathrm{S^{t}}=\left\{s_{1}^{t},s_{2}^{t},s_{3}^{t}\ldots s_{k}^{t},acc_{1}^{t},acc_{2}^{t},acc_{3}^{t}\ldots acc_{k}^{t},acc_{global}^{t}\right\}$ (16)
Action: In each stage, we need to reassign each model’s weights based on the model’s quality. The action space is shown as 17. $a_{i}^{t}$ denotes the weight value assigned to the i-th client in the state t, while the sum of the corresponding weight values of all clients is 1.
$\mathrm{A^{t}}=\left\{a_{1}^{t},a_{2}^{t},...a_{k}^{t}\right\},{\textstyle\sum_{1}^{k}}a_{i}^{t}=1,a_{i}^{t}\in(0,1)$ (17)
Reward: We define the model accuracy aggregated using the weighted average method as $Acc_{all}$. Also, we define the global model accuracy obtained by the overall model weights assigned to the client for the mth time as $Acc_{m}$. We calculate the reward value by subtracting the difference of $Acc_{m}$ from $Acc_{all}$. If the calculated result is greater than zero, this indicates that the weights assigned by the agent improve the accuracy of the global model, and we give a positive reward. Conversely, we give a penalty reward. $\varphi$, $\phi$ denotes the reward and penalty factors, respectively. So the reward is defined as equation 18.
$Reward=\begin{cases}\varphi*(Acc_{m}-Acc_{all}),Acc_{m}>Acc_{all}\\ \phi*(Acc_{m}-Acc_{all}),Acc_{m}<Acc_{all}\end{cases}$ (18)
We use multiple parallel environments to accelerate the sampling to speed up the training agent in this phase. The stage is shown in figure 1. This stage includes a central Learner and multiple Worker nodes. Each Worker corresponds to a parallel environment. The workflow of each Worker is as follows: first, each Worker performs global model fusion based on the assigned weights; then verifies the accuracy of the global model by interacting with the parallel environment; and finally receives the reward values from the parallel environment feedback. Finally, each Worker stores the corresponding ones in the sampling buffer pool. Multiple workers interact with each environment independently, thus achieving parallel sampling to improve the sampling efficiency. After each worker collects a certain batch of samples, Learner trains the agent by taking a certain amount of sample data from the experience pool.
We use the TD3 algorithm for this phase of the method design. The TD3 algorithm consists of six network models, including an Actor-network $P(w)$, two Critic networks $Q_{1}\left(\theta_{1}\right),Q_{2}\left(\theta_{2}\right)$, and a target Actor-network $P^{\prime}(w)$, two target Critic networks $Q^{\prime}_{1}\left(\theta_{1}\right),Q^{\prime}_{2}\left(\theta_{2}\right)$. The network results of the TD3 algorithm are shown in figure 1. The learner randomly draws N batches of sample data from the buffer pool every certain round for training the model. The training processes are as follows. The process is shown in algorithm 2
[…]
(2) The TD target value is calculated using equation 19, where $\operatorname{Min}\left(q_{1},q_{2}\right)$ takes the minimum value of both.
$y_{t}\leftarrow r+\gamma\operatorname{Min}(q1,q2)$ (19)
(3) Select the action based on the Actor-network, input the state and action into the Critic network separately, and let these two networks output the corresponding prediction reward sum.
(4) Calculate the TD error. The calculation formula is as equation 20.
$\delta_{1,t}=q_{1,t}-y_{t},\quad\delta_{2,t}=q_{2,t}-y_{t}$ (20)
(5) Update the Critic network as equation 21.
$\begin{array}[]{l}\theta_{1}\leftarrow\theta_{1}-\alpha\cdot\delta_{1,t}\cdot\nabla_{w}Q_{1}\left(s_{t},a_{t};\theta_{1}\right)\\ \theta_{2}\leftarrow\theta_{2}-\alpha\cdot\delta_{2,t}\cdot\nabla_{w}Q_{2}\left(s_{t},a_{t};\theta_{2}\right)\end{array}$ (21)
(6) Update the strategy network every d rounds through the Actor-network output action as equation 22.
$\mathrm{w}\leftarrow\mathrm{w}+\beta\cdot\nabla_{w}P\left(s_{t};\mathrm{w}\right)\cdot\nabla_{w}Q_{1}\left(s_{t},a_{t};\theta_{1}\right)$ (22)
(7) Update the target Actor and Critic network parameters every d rounds as equation 23.
$\begin{array}[]{l}w^{\prime}\leftarrow\tau w+(1-\tau)w^{\prime}\\ \theta_{1}^{\prime}\leftarrow\tau\theta_{1}+(1-\tau)\theta_{1}^{\prime}\\ \theta_{2}^{\prime}\leftarrow\tau\theta_{2}+(1-\tau)\theta_{2}^{\prime}\end{array}$ (23)
Algorithm 2 FedDRL Stage 2 Algorithm Process
1: Client Models $\left\{\theta_{1}^{t},\theta_{2}^{t},\theta_{3}^{t},\ldots\theta_{\mathrm{n}}^{t}\right\}$, Round R, Worker Number N, Buffer Memory Pool M Initialize Learner Parameters: Actor Parameter P$\left(w\right)$, Critic Network $Q_{1}\left(\theta_{1}\right),Q_{2}\left(\theta_{2}\right)$ Target Actor Parameter $P^{\prime}\left(w^{\prime}\right)$, Target Critic Network $Q_{1}^{\prime}\left(\theta_{1}^{\prime}\right),Q_{2}^{\prime}\left(\theta_{2}^{\prime}\right)$ $w^{\prime}\leftarrow w,\theta_{1}^{\prime}\leftarrow\theta_{1},\theta_{2}^{\prime}\leftarrow\theta_{2}$
2: Optimized Client Model Weight $W=\left\{w_{1}^{t},w_{2}^{t},\ldots w_{\mathrm{k}}^{t}\right\}$
3:/* Each Worker Sampling Step */
4:for $worker$ from 1 to $N$ do
5:     $a_{t}\leftarrow P\left(s_{t},w\right)$ // Randomly Choose an Act from $P\left(s_{t},w\right)$
6:     $\left\{w_{1}^{t},w_{2}^{t},w_{3}^{t}\ldots w_{k}^{t}\right\}\leftarrow Step\left(a_{t}\right)$
7:     $\theta_{global}^{t}\leftarrow Agg\left(\sum_{i=1}^{k}w_{i}^{t}\theta_{i}\right)$
8:     $R_{t}\leftarrow CaculateReward\left(ACC_{t}-ACC_{avg}\right)$
9:     $M\leftarrow Store\left(<S_{t},A_{t},R_{t},S_{t+1}>\right)$
10:/* Center Learner Training Step */
11:for $r$ from 1 to R do
12:     Randomly Sampling N Batches of Data from M
13:     $a_{t+1}^{\prime}\leftarrow P^{\prime}\left(s_{t+1}\right)$
14:     $y\leftarrow r+\gamma Min\left(Q_{1}^{\prime}\left(s_{t+1},a_{t}^{\prime}\right),Q_{2}^{\prime}\left(s_{t+1},a_{t}^{\prime}\right)\right)$
15:     Update Critic Network $\theta_{1}\leftarrow argmin_{\theta_{1}}\frac{1}{N}\sum\left(y-Q_{\theta_{1}}\left(s,a\right)\right)^{2}$
16:     Update Critic Network $\theta_{2}\leftarrow argmin_{\theta_{2}}\frac{1}{N}\sum\left(y-Q_{\theta_{2}}\left(s,a\right)\right)^{2}$
17:     Every d Rounds:
18:     Update Actor Network: $\nabla_{w}J\left(w\right)=N^{-1}\sum\nabla_{w}Q_{\theta_{1}}\left(s,a\right)\mid_{a=P\left({s}\right)}\nabla_{w}P\left({s}\right)$
19:     Update Target Critic Network: $\theta_{1}^{\prime}\leftarrow\tau\theta_{1}+\left(1-\tau\right)\theta_{1}^{\prime},\theta_{2}^{\prime}\leftarrow\tau\theta_{2}+\left(1-\tau\right)\theta_{2}^{\prime}$
20:     Update Target Actor Network: $w^{\prime}\leftarrow\tau w+\left(1-\tau\right)w^{\prime}$
21:After R Rounds, Save Trained Model
22:Output Optimized Model Weight W = $\left\{w_{1}^{t},w_{2}^{t}\ldots w_{k}^{t}\right\}$
4 SYSTEM DESIGN
We propose a trusted federated model fusion framework(FedDRL) to solve the client weight assignment problem. The main workflow is divided into five steps. Figure 2 shows the system architecture of FedDRL.
Figure 2: The system architecture of FedDRL
Step 1: First, each client downloads the global model, initializes the model parameters using the global model, and then trains the model using local private data.
Step 2: After each client’s training, each client uploads its respective model parameters to the server side.
Step 3: After receiving the client’s model parameters, the server side performs the trusted client selection and adaptive weight assignment in two stages according to the FedDRL algorithm.
In stage 1, the server uses the SelectTrustClient(.) algorithm to train the agent, then updates each network parameter of the agent, and finally outputs the optimal action to achieve the selection of trusted clients.
In stage 2, the server uses the AdaptCalculateWeight(.) algorithm to train the agent, then updates each network parameter of the agent, and finally achieves the optimal weight assignment for each client through the agent.
Step 4: The server side fuses the client models based on the calculated optimal weights and finally fuses the global models.
Step 5: The server side distributes the global model to each client.
We set several communication rounds and repeat steps 1-5 until the global model converges. The process is shown in Algorithm 3.
Algorithm 3 The FedDRL framework
1: Private Dataset $\left\{D_{1},D_{2},\ldots D_{\mathrm{n}}\right\}$, communication round E
2: The Global model $\left\{M_{global}\right\}$
3:/* Client Process */
4:for $C_{i}$ from 1 to $N$ do
5:     $M_{i}\leftarrow\operatorname{GetGlobalModel}(round=i)$ // Get global model and init cleint model
6:     $M_{i}\leftarrow\operatorname{TrainLocalModel}(D_{i})$ // Train model $M_{i}$ based Dataset $\left\{D_{i}\right\}$
[…]
8:/* Server Process */
9:for $e$ from 1 to $E$ do
10:      Store ($(\left\{M_{1},M_{2},\ldots M_{n}\right\})\leftarrow\operatorname{Receive}\left(M_{i}\right)$ // Receive Client Model /* FedDRL Algorithm Process */
11:     Train the Stage 1 Agent
12:     Update the $\operatorname{SelectTrustClient}\left(.\right)$ Algorithm parameters // According Algorithm 1
13:     $\left\{M_{a},M_{b},\ldots M_{\mathrm{k}}\right\}\leftarrow\operatorname{SelectTrustClient}\left(\left\{M_{1},M_{2},\ldots M_{n}\right\}\right)$
14:     Train the Stage 2 Agent
15:     Update the $\operatorname{AdaptCalculateWeight}\left(.\right)$ Algorithm parameters // According Algorithm 2
16:     $\left\{W_{1},W_{2},\ldots W_{\mathrm{n}}\right\}\leftarrow\operatorname{AdaptCalculateWeight}\left(\left\{M_{a},M_{b},\ldots M_{k}\right\}\right)$
17:     $M_{global}\leftarrow\operatorname{FusionGlobalModel}\left(\left\{W_{1},W_{2},\ldots W_{n}\right\}\right)$
18:     $C_{i}\leftarrow\operatorname{SendGlobalModel}\left(M_{global}\right)$
5 EXPERIMENT
5.1 Experiment setup
5.1.1 Experiment datasets
We validate the FedDRL framework on three image classification datasets.
Fashion-MNIST dataset: Fashion-MNIST consists of 70,000 samples in 10 classes.
CIFAR-10 dataset: CIFAR-10 dataset consists of 60000 images in 10 classes.
CIFAR-100 dataset: CIFAR-100 has 100 classes containing 600 images for each class.
Data set partitioning: To simulate the distribution of non-IID datasets on the clients, we use the method in the paper [26] to partition different clients’ data on the open-source dataset using the Dirichlet function. The Dirichlet function could simulate the data partitioning of each client by setting the alpha parameter. When the alpha parameter tends to zero, the data distribution of clients is assigned to only a certain class of the public dataset; when the alpha parameter tends to positive infinity, the closer the distribution of client data, which is the amount and class of data is consistent across clients. To simulate the distribution of non-iid data, we set alpha to 1 to split the three datasets into ten clients. The data distribution is shown in Figure 3.
Refer to caption (a) fashionmnist
Refer to caption (b) CIFAR-10
Refer to caption (c) CIFAR-100
Figure 3: The non-iid distribution of 10 clients(alpha=1).
5.1.2 Comparison of methods
We compared the FedDRL algorithm with two traditional federated learning algorithms.
FedAvg[1]: The FedAvg method is the baseline of federated learning, which allocates the weight of the client model by calculating the ratio of the number of samples from each client to the total number of samples.
FedProx[2]: this algorithm improves the FedAvg method by adding a regularization term to the client model to optimize the performance of the federated learning.
5.1.3 Experimental metrics
[…]
$A_{\text{gobal }}=\left\{A_{\text{gobal }}^{(1)}\ldots A_{\text{gobal }}^{(t)}\ldots A_{\text{gobal }}^{(m)}\right\}$ (24)
To demonstrate the behavior of each client. We also calculated the accuracy of the local model for each client in each round, which is expressed as 25.
$A_{\text{c}}=\left\{A_{\text{c}}^{(1)}\ldots A_{\text{c}}^{(t)}\ldots A_{\text{c}}^{(m)}\right\},c\in\left[1,2...n\right]$ (25)
5.1.4 Experimental configuration
Hardware configuration: We completed the experiments using a high-performance workstation with an Intel i9-12900k processor, 64GB RAM, and an NVIDIA RTX3090 graphics card.
Software configuration: We used two frameworks to complete the Federated Learning and Reinforcement Learning experiments. We use the FedBolt, our self-developed experimental framework, to conduct the Federated Learning experiments. We use this framework to simulate a Federated Learning experiment environment with a different number of clients and different data distributions. For the training process of the reinforcement learning model, we use the stablebaseline3 framework [27], based on which we design two algorithms for trusted client selection and model weight assignment.
Federeted Learning Network setup: We designed different network models for different datasets. For the CIFAR-10 and CIFAR-100 datasets, we create a 6-layer CNN network to train the model. We designed a 4-layer MLP network for the FASHION-MNIST dataset.
DRL Network setup: Since we use a staged reinforcement learning strategy, we use a distributed A2C algorithm in the first stage. Each worker and the central node include a 6-layer MLP for the Actor and a 6-layer MLP for the Critic Network. In the second stage, we used the TD3 algorithm and a 6-layer MLP network for each module of TD3.
5.2 Experimental Results
We validate the FedDRL algorithm in three experiments. These include client attack experiments, low-quality model fusion experiments, and hybrid experiments. The client attack experiments focus on validating our proposed trusted client selection algorithm of stage 1. The low-quality model experiment mainly validates our proposed adaptive weight calculation algorithm of stage 2. The hybrid experiment contains both client attack and low-quality model components and mainly verifies the overall performance of the FedDRL method.
5.2.1 Malicious Client Attack Experiment
We define three attack types to simulate the client-side attack behavior in federated learning. To accomplish the experiments, we simulated different numbers of clients and different types of attacks in three datasets. The experimental configuration of the client attack is shown in table 1.
Type 1: The client directly uploads the initialized model or makes the model accuracy less than 10% by modifying the model’s hyperparameters.
Type 2: We use falsified data to perform the attack. We use a certain percentage of forged data to participate in model training (e.g., mix the cifar10 dataset with $80\%$ of cifar100 data and generate these cifar100 data labels as cifar10 corresponding label types). We conduct the attack by faking sample data to train the client’s local model, thus reducing the client model’s accuracy.
Type 3: We select some clients to simulate the attack and divide the training process of these clients into standard and attack rounds. In the standard round, each client does not perform the attack behavior. Instead, each client deliberately uploads the prepared malicious model in the attack round. We also set that these clients alternately initiate the attack behavior.
Table 1: Client Attack Experimental Settings for Different Numbers of Clients
Number Attack Type Malicious ID Number of samples Accuracy of models ($\leq$)
Client=5 Type 1 Client1 7750 $A\leq 10\%$
Type 2 Client1 7750 $10\%\leq A\leq 20\%$
Type 3 Client1 7750 Attack round $A\leq 10\%$
Client=10 Type 1 Client1,Client6 4222, 4938 $A\leq 10\%$
Type 2 Client1,Client6 4222, 4938 $10\%\leq A\leq 20\%$
Type 3 Client1,Client6 4222, 4938 Attack round $A\leq 10\%$
Client=15 Type 1 Client1,Client6,Client11 3670, 3314, 4454 $A\leq 10\%$
Type 2 Client1,Client6,Client11 3670, 3314, 4453 $10\%\leq A\leq 20\%$
Type 3 Client1,Client6,Client11 3670, 3314, 4453 Attack round $A\leq 15\%$
According to the experimental setup, we compared the FedDRL algorithm with the two baseline algorithms. In the attack experiments, we set the total number of communication rounds to 100 rounds, and each node performs local model training with 1 epoch. To show the attack behavior of each client and the accuracy of different algorithms in each attack scenario in more detail, we counted the accuracy of each client’s local model and the accuracy of the server-side global model in each communication round. The specific experimental results are shown in table 2.
Table 2: Global model accuracy for different number of client attack types
DataSet Method Clients=5 Clients=10 Clients=15
Type1 Type2 Type3 Type1 Type2 Type3 Type1 Type2 Type3
fashion -mnist FedAvg 0.862 0.875 0.863 0.792 0.881 0.821 0.776 0.878 0.812
FedProx 0.873 0.884 0.864 0.791 0.882 0.824 0.764 0.879 0.811
Ours 0.885 0.878 0.883 0.877 0.886 0.887 0.881 0.886 0.882
Cifar10 FedAvg 0.596 0.691 0.751 0.335 0.681 0.314 0.139 0.664 0.197
FedProx 0.586 0.732 0.775 0.331 0.716 0.363 0.122 0.719 0.202
Ours 0.731 0.701 0.747 0.694 0.711 0.727 0.679 0.702 0.689
Cifar100 FedAvg 0.376 0.412 0.298 0.273 0.416 0.287 0.162 0.398 0.176
FedProx 0.398 0.442 0.321 0.223 0.436 0.208 0.172 0.426 0.183
Ours 0.412 0.438 0.431 0.426 0.432 0.421 0.423 0.412 0.422
To show the effect of the FedDRL algorithm on global model fusion at each communication round. We conducted experiments using the CIFAR10 dataset on 5, 10, and 15 clients. And We compared FedDRL with FedAvg and FedProx algorithms for the global model accuracy.
We analyze the experimental results for different numbers of client models and different client data. In attack type 1, our algorithm outperforms the FedAvg and FedProx algorithms on all three data sets. In malicious data attack type 2, our algorithm outperforms the FedAvg algorithm and slightly underperforms the FedProx algorithm alone. In the attack type 3 scenario, our algorithm outperforms the comparison algorithm in most cases, especially when multiple malicious clients are involved in model fusion. To show the relationship between the global and client model’s accuracy in each attack scenario. We conducted more detailed experiments on the Cifar10 dataset. The experimental results are shown in Figure 4, Figure 5, and Figure 6.
In attack type 1, our trained agent can dynamically select trusted clients for model fusion and eliminate malicious models from participating, so our algorithm has higher reliability.
In attack type 2, our algorithm is lower than FedProx because the FedProx algorithm uses control parameters to force the models of all client nodes to converge to the global model, which will improve the accuracy of the global model by improving the accuracy of the malicious model to some extent. Our trained reinforcement learning agent will filter out low-accuracy models to participate in the fusion after several communication rounds.
In attack type 3, the trained agent can dynamically adjust the trusted clients, so our algorithm works reliably even in the case of rotating attacks by malicious clients.
Refer to caption (a) 5 clients type1 FedAvg
Refer to caption (b) 5 clients type1 FedProx
Refer to caption (c) 5 clients type1 FedDRL
Refer to caption (d) 10 clients type1 FedAvg
Refer to caption (e) 10 clients type1 FedProx
Refer to caption (f) 10 clients type1 FedDRL
Refer to caption (h) 15 clients type1 FedAvg
Refer to caption (i) 15 clients type1 FedProx
Refer to caption (j) 15 clients type1 FedDRL
Figure 4: Global model accuracy in client attack type 1
Refer to caption (a) 5 clients type3 FedAvg
Refer to caption (b) 5 clients type3 FedProx
Refer to caption (c) 5 clients type3 FedDRL
Refer to caption (d) 10 clients type3 FedAvg
Refer to caption (e) 10 clients type3 FedProx
Refer to caption (f) 10 clients type3 FedDRL
Refer to caption (h) 15 clients type3 FedAvg
Refer to caption (i) 15 clients type3 FedProx
Refer to caption (j) 15 clients type3 FedDRL
Figure 5: Global model accuracy in client attack type 2
Refer to caption (a) 5 clients type3 FedAvg
Refer to caption (b) 5 clients type3 FedProx
Refer to caption (c) 5 clients type3 FedDRL
Refer to caption (d) 10 clients type3 FedAvg
Refer to caption (e) 10 clients type3 FedProx
Refer to caption (f) 10 clients type3 FedDRL
Refer to caption (h) 15 clients type3 FedAvg
Refer to caption (i) 15 clients type3 FedProx
Refer to caption (j) 15 clients type3 FedDRL
Figure 6: Global model accuracy in client attack type 3
5.2.2 Low-quality model fusion experiments
To validate the performance of our proposed algorithm FedDRL, we performed validation on three datasets, Fashion-MNIST, CIFAR-10, and CIFAR-100. Since all three are open source, these data are high-quality data. Therefore each client used these sample data to train their respective models, and these models did not exhibit significant accuracy differences. We set up the following experiments to verify the impact of low-quality models involved in global model fusion in a real-world scenario.
(1) To verify the impact of different numbers of customer nodes on the algorithm, we conducted experiments on three datasets with 5, 10, and 15 clients, respectively. We also set the parameter alpha=1 of Dirichlet to complete the dataset partitioning corresponding to each client.
(2) We introduced a method for low-quality models to participate in the fusion. At the beginning of each round of federated learning, these clients initialize the local model using the global model. After several rounds of communication, when the accuracy of the local model reaches a certain accuracy, we simulate the participation of low-quality models in global model fusion by controlling the accuracy of these client models (e.g., model accuracy is kept between 50% and 60%). These clients upload low-quality models to the server side. Our experimental configuration of low-quality simulation for a different number of clients in table 3.
Table 3: Experimental Settings for Low-quality Model experiments
Number Dataset Low-quality Model ID Number of samples Accuracy of models ($\leq$)
Client=5 Fashion-MINST Client1 9061 53%
CIFAR-10 Client1 7750 52%
CIFAR-100 Client1 9278 22%
Client=10 Fashion-MINST Client1,Client5 5071, 7245 51%, 52%
CIFAR-10 Client1,Client5 4222, 6039 50%, 54%
CIFAR-100 Client1,Client5 4191, 5491 22%
Client=15 Fashion-MINST Client1,Client5,Client10 4405, 3752, 1809 52%, 51%, 53%
CIFAR-10 Client1,Client5,Client10 3670, 3128, 1509 49%, 52%, 55%
CIFAR-100 Client1,Client5,Client10 3073, 3494, 2910 22% 19% 23%
We compared the FedDRL algorithm with both FedAvg and FedProx algorithms. We set the number of communication rounds to 100. Each client performs one epoch of local model training. The experimental results are shown in table 4.
Table 4: Global model accuracy of Low-quality model experiments
Method Fashion-MINST CIFAR-10 CIFAR-100
Clients=5 Clients=10 Clients=15 Clients=5 Clients=10 Clients=15 Clients=5 Clients=10 Clients=15
FedAvg 0.857 0.858 0.841 0.705 0.664 0.602 0.386 0.373 0.365
FedProx 0.865 0.861 0.829 0.714 0.652 0.607 0.402 0.391 0.386
Ours 0.885 0.887 0.884 0.725 0.706 0.698 0.422 0.418 0.407
To demonstrate the FedAvg, FedProx, and FedDRL algorithms for global model fusion at each communication round. We have taken the CIFAR10 dataset as an example and conducted experimental comparisons with 5, 10, and 15 clients, respectively. The experimental results of each algorithm are shown in Figure 5.
The comparison experiments of CIFAR-10 show that the performance of FedAvg and FedProx algorithms keeps decreasing as the number of low-quality models involved in fusion increases. Also, these two algorithms fluctuate up and down with the accuracy of the low-quality models due to the number of samples used by these two algorithms to calculate the global model. Therefore, when the weight value of the model is fixed, the change in the accuracy of the model affects the change in the accuracy of the global model. Meanwhile, FedDRL has a faster convergence speed than both FedAvg and FedProx algorithms because the FedDRL algorithm can adjust the corresponding weights adaptively according to the quality of each client model, which mitigates the impact of low-quality models on the accuracy of the global model and thus accelerates the convergence speed of the global model to a certain extent.
Refer to caption (a) 5 clients FedAvg
Refer to caption (b) 5 clients type2 FedProx
Refer to caption (c) 5 clients FedDRL
Refer to caption (d) 10 clients FedAvg
Refer to caption (e) 10 clients FedProx
Refer to caption (f) 10 clients FedDRL
Refer to caption (g) 15 clients FedAvg
Refer to caption (h) 15 clients FedProx
Refer to caption (i) 15 clients FedDRL
Figure 7: Global model accuracy in Low-quality model.
5.2.3 Hybrid experiment
In this section, we fuse two attack types (type 1 and type 3) and simulate a low-quality model to participate in the fusion. We verify The performance of the FedDRL algorithm in this hybrid Scenario. We also compare with FedAvg and FedProx methods.
We used the cifar10 dataset and conducted experiments with 10 and 15 client cases. Client 1 only uploads the initialization model for each training, Client 6 simulates low-quality models to participate in fusion, client 10 or Client 11 initiates attack behavior in odd-numbered communication rounds, and even-numbered rounds participate in training normally. The remaining nodes normally participate in each round of Federated learning tasks. The detailed configuration of the experiments is shown in table 5. After 100 communication rounds, the global model accuracy of each algorithm is shown in table 6.
Table 5: Hybrid Experimental Settings for Different Numbers of Clients
Number Client ID Type Number of samples Model Accuracy
Client=10 Client1 Attack Type 1 4222 $A\leq 10\%$
Client6 Low-quality Model 4938 $45\%\leq A\leq 50\%$
Client10 Attack Type 3 3560 Attack round $A\leq 15\%$
Client=15 Client1 Attack Type 1 3670 $A\leq 10\%$
Client6 Low-quality Model 3314 $45\%\leq A\leq 50\%$
Client11 Attack Type 3 4453 Attack round $A\leq 15\%$
Table 6: Global model Accuracy of Hybrid Experimental in different clients’ number
Method Fashion-MINST CIFAR-10 CIFAR-100
Clients=10 Clients=15 Clients=10 Clients=15 Clients=10 Clients=15
FedAvg 0.835 0.823 0.368 0.348 0.223 0.238
FedProx 0.821 0.846 0.308 0.341 0.241 0.266
Ours 0.876 0.883 0.701 0.698 0.426 0.418
Refer to caption (a) 10 clients FedAvg
Refer to caption (b) 10 clients FedProx
Refer to caption (c) 15 clients FedDRL
Refer to caption (d) 15 clients FedAvg
Refer to caption (e) 15 clients FedProx
Refer to caption (f) 15 clients FedDRL
Figure 8: Comparison of global model accuracy between different algorithms.
The hybrid scenario experiments show that the FedAvg and FedProx algorithms can no longer perform global model fusion properly due to the attack behavior of the client. Meanwhile, malicious models’ involvement in the global model fusion makes the accuracy of the global model based on the traditional algorithm very low. In the FedDRL algorithm, we adopt the two-stage method. In the first stage, we eliminate malicious models from participating in a fusion. We use an adaptive weight fusion strategy in the second stage, reducing low-quality models’ influence. Thus our algorithms work correctly even in this hybrid scenario. The global model and the model accuracy of each client for each communication round of these three algorithms are shown in Figure 8.
5.2.4 Discussion
We analyze the experimental results for different model fusion scenarios as well as for different numbers of client models. In the client-side attack experiments, under attack type 1, our algorithm works better than the FedAvg method on all three datasets. Meanwhile, as the number of malicious nodes increases, the global model accuracy of both FedAvg and FedProx algorithms shows a serious decrease. In contrast, our algorithm still maintains a smooth convergence trend. Under attack type 2, our algorithm is better than the FedAvg algorithm and slightly lower than the FedProx algorithm. Under attack type 3, our algorithm outperforms the edAvg and FedProx algorithms on all three datasets, which no longer work under rotating attacks with multiple malicious clients. In contrast, our algorithm can work reliably. In the low-quality model experiments, our algorithm outperforms the FedAvg and FedProx algorithms on all three datasets. In the hybrid experiments, the FedAvg and FedProx algorithms no longer work properly, and our algorithm can fuse the global models properly.
Since the FedAvg and FedProx methods use the number of samples to calculate the model weights, they cannot adaptively adjust the responses of malicious and low-quality models. They thus cannot be applied to these attack scenarios. Our proposed method is based on staged reinforcement learning for malicious node removal and optimal weight assignment for trusted clients and thus can work reliably in various environments.
6 CONCLUSION
To solve the problem of assigning client weights due to data heterogeneity in a federated learning environment. We propose FedDRL, an adaptive weight calculation method based on staged reinforcement learning. The method comprises two stages: selecting trusted clients and adaptive weight assignment. In the first stage, our approach can eliminate malicious client models from participating in the global model fusion. It also can dynamically select trusted client models to participate in the next stage of model weight assignment. In the second stage, Our method can adaptively calculate the corresponding weights based on the model quality of each client. Also, we propose a distributed reinforcement learning method to accelerate the agent’s training. Finally, we proposed five types of model fusion scenarios to validate our approach, and experiments show that our proposed algorithm can work reliably in all kinds of model fusion scenarios while maintaining the accuracy of the global model.
Currently, our algorithm requires separate training of the agent according to the client fusion scenarios. In future work, we will extend the proposed algorithm to accommodate model fusion tasks for different task scenarios. We will also continue improving the algorithm’s performance to achieve more client tasks.

### 19 Mar 2024 v4

@@
 Abstract
-Traditional federated learning uses the number of samples to calculate the weights of each client model and uses this fixed weight value to fusion the global model.
-However, in practical scenarios, each client’s device and data heterogeneity leads to differences in the quality of each client’s model.
-Thus the contribution to the global model is not wholly determined by the sample size.
-In addition, if clients intentionally upload low-quality or malicious models, using these models for aggregation will lead to a severe decrease in global model accuracy.
-Traditional federated learning algorithms do not address these issues.
-To solve this probelm, we propose FedDRL, a model fusion approach using reinforcement learning based on a two staged approach.
-In the first stage, Our method could filter out malicious models and selects trusted client models to participate in the model fusion.
-In the second stage, the FedDRL algorithm adaptively adjusts the weights of the trusted client models and aggregates the optimal global model.
-We also define five model fusion scenarios and compare our method with two baseline algorithms in those scenarios.
-The experimental results show that our algorithm has higher reliability than other algorithms while maintaining accuracy.
+Federated learning facilitates collaborative data analysis among multiple participants while preserving user privacy.
+However, conventional federated learning approaches, typically employing weighted average techniques for model fusion, confront two significant challenges: (1) The inclusion of malicious models in the fusion process can drastically undermine the accuracy of the aggregated global model.
+(2) Due to the heterogeneity problem of devices and data, the number of client samples does not determine the weight value of the model.
+To solve those challenge, we propose a trustworthy model fusion method based on reinforcement learning (FedDRL), which includes two stages.
+In the first stage, we propose a reliable client selection mechanism to exclude malicious models from the fusion process.
+In the second stage, we propose an adaptive model fusion method that dynamically assigns weights based on model quality to aggregate the best global models.
+Finally, We validate our approach against five distinct model fusion scenarios, demonstrating that our algorithm significantly enhances reliability without compromising accuracy.
 1 Introduction
-With the development of deep learning technologies, numerous industries are applying them to their respective fields, resulting in the emergence of intelligent transportation, smart logistics, and smart healthcare models.
-These technologies reduce the industry’s production and management costs, increase operational efficiency, and accelerate the industry’s digitization process.
-However, most current deep learning-based model training uses a supervised learning method.
-The number and diversity of samples is the key to training a high-quality model.
-Therefore, how to obtain massive and diverse data samples has become the first step to training deep learning models.
-This method leads to the source of these samples being no longer limited to a single industry but to multiple industries collaborating to construct large-scale datasets.
-Moreover, with the emergence of cross-industry collaborative data analysis models, data analysis sources are no longer limited to a single organization but involve multiple parties.
-Although the participation of multiple parties in the data analysis process can solve the problem of cross-domain data analysis, this strategy also faces a new challenge.
-The traditional data analysis method requires uploading each participant’s raw data to a central node and training models on the central node.
-However, this workflow involving raw data transfer poses security and privacy risks.
-Meanwhile, raw data transmission is not permitted for data-sensitive industries, such as hospital patient and financial industry data.
-How to achieve federated data analysis of multiple parties while protecting data security is becoming hot research.
-Federated learning is an emerging paradigm for distributed computing.
-The structure of federated learning has a central node and multiple participants, where multiple participants train models with their private data, and the central node achieves knowledge extraction and information fusion of each participant by fusing the models uploaded by each client.
-Since the process does not directly use the participants’ data, federated learning can achieve joint data analysis of multiple participants while protecting user privacy in advance.
-To solve mobile device word input privacy protection issues, Google proposed federated learning and the FedAvg method [1], which ensure user privacy by transmitting models instead of raw data.
-Increasingly, companies are applying federated learning in various domains to tackle cross-industry data analysis challenges.
-However, in practical federated learning environments, the heterogeneity of data among participants gives rise to the non-iid problem, which makes model convergence challenging during the fusion stage of the FedAvg method.
-Additionally, variations in computational resources among participants lead to slow convergence of the global model.
-To address these challenges, some research has optimized the global model convergence problem by adding control parameters in the client-side model fusion process, such as FedProx [2], Scaffold [3], and FedNova [4].
-The Moon algorithm [5] optimizes the convergence of the global model by introducing a comparative loss between the client-side model and the global model.
-There is also some research [6] [7] to improve the accuracy of the global model by selecting the clients involved in the global model.
-These algorithms somewhat solve the global model’s convergence problem in the non-iid data among participants.
-Current traditional federal learning methods, such as FedAvg and FedProx, use a ratio based on the number of samples of each participant to the total number of samples to calculate the model weights of each participant.
-However, in a practical scenario, the method has two problems.
-On the one hand: the method does not consider the different participants’ model quality, which gives higher weights to the client with a larger number of samples.
-However, due to the variability of computational resources and sample quality across clients, it does not mean that a client can train a higher-quality model when a client has more samples.
-Therefore, using these algorithms assumes that the sample quality and computational resources are consistent across clients.
-On the other hand, if the participants have malicious attack behavior and intentionally upload malicious models to participate in fusion, which will seriously influence the global model convergence process.
-Therefore it becomes increasingly important to design adaptive defenses for federated learning systems [8].
-How to eliminate malicious models in federated learning model fusion has become a vital issue.
-At the same time, when the client uploads low-quality models to participate in fusion, how to adaptively adjust each model’s fusion weights based on the models’ quality is also an urgent problem to be solved.
-Some research has used reinforcement learning methods to address these issues of weighting.
-Such as, Favor [7] method uses the DDPG to calculate the weights of the participant models.
-Some research has also used reinforcement learning to solve the problems of device selection [9] [10], resource optimization [11] [12], and communication optimization in IoT federated learning scenarios.
-Reinforcement learning (RL) methods use a try-and-learn strategy.
-The core of the reinforcement learning approach is to train an intelligent agent.
-The agent uses different actions to interact with the external environment.
-The environment provides reward and punishment values based on the selected actions of the intelligence.
-The intelligent body achieves optimal action selection by maximizing the reward value.
-However, when we use reinforcement learning, we must address some issues.
-On the one hand, we must collect training samples by continuously interacting with the environment.
-When the cost of interacting with the external environment is too high, or the interaction cost is unacceptable (e.g., driverless scenario), how to collect samples quickly will directly affect the training time of reinforcement learning.
-On the other hand, when the action space of the agent is high-latitude and continuous, this will lead to an increasing sampling time.
-Therefore, due to these two problems, traditional single-agent reinforcement learning training methods can take much time.
-When we apply reinforcement learning to federated learning, we must solve those problems.
-As the number of participants grows, the time cost of training agents will increase.
-How to optimize the action space for reinforcement learning and speed up the training process of the agent is an urgent problem to be solved.
-To address these issues, we propose a staged reinforcement learning algorithm (FedDRL) to solve the problem of model weight calculation in federated learning.
-The algorithm consists of two phases: the selection of trusted clients and the adaptive assignment of weights.
-We use a staged approach to accelerate the agent’s training.
-We use a distributed reinforcement learning algorithm to speed up the collection of training samples.
-The contributions of this paper is as follows.
-  • We propose a federated learning model fusion method based on reinforcement learning, which solves the problem of malicious clients’ participation in model fusion and clients’ model weight assignment.
-  • We propose a trusted client selection strategy that eliminates malicious models from participating in global model fusion while enabling the selection of trusted clients.
-  • We propose a adaptive weight assignment method that adaptively adjusts the weights corresponding to the clients according to the quality of their uploaded models and therefore computes the optimal global model.
+With the advent of deep learning technologies, various industries have been integrating these technologies into their sectors, promoting the development of intelligent transportation, smart logistics, and healthcare systems.
+These technologies are crucial in reducing production and management costs, enhancing operational efficiency, and accelerating industry digitization.
+However, supervised learning remains the primary method for training deep learning models, where the volume and diversity of samples are essential for creating high-quality models.
+Consequently, acquiring extensive and varied data samples has emerged as the initial step in training deep learning models.
+This approach has led to sample sources expanding from single industries to collaborations across multiple sectors to develop large-scale datasets.
+To achieve multi-party joint data analysis under the condition of protecting data security and privacy, Google has proposed federated learning technology for the first time.
+Although federated learning solves the problem of user privacy protection, the traditional federated learning algorithm assumes that all participants are trustworthy.
+On the contrary, in the actual scenario, if participants exhibit malicious behavior and intentionally contribute harmful models to the fusion process, it can significantly disrupt the global model’s convergence.
+Thus, creating adaptive defenses for federated learning systems becomes increasingly crucial [9].
+Identifying methods to remove malicious models in federated learning model fusion has become a critical issue.
+Simultaneously, when a client submits low-quality models for fusion, determining how to adaptively adjust each model’s fusion weights based on their quality is also an urgent problem needing resolution.
+Some studies have applied reinforcement learning techniques to address these weighting issues.
+For instance, the Favor [10] method uses the DDPG to assign weights to participant models.
+Additional research has applied reinforcement learning to address device selection [11] [12], resource optimization [13] [14], and communication optimization in IoT federated learning contexts.
+Reinforcement learning (RL) employs a trial-and-error strategy.
+The essence of this approach is training an intelligent agent that interacts with the external environment through varied actions.
+The environment then provides feedback in the form of rewards and penalties based on the agent’s actions, guiding the agent toward optimal action selection by maximizing reward value.
+However, employing reinforcement learning presents certain challenges.
+Firstly, continuous training is required for sample collection through environmental interaction.
+When the cost of such interactions is prohibitive or unacceptable (for example, in our scenario, where the server must frequently calculate the global model’s parameters), the efficiency of sample collection significantly impacts the reinforcement learning training duration.
+Secondly, when the agent’s action space is vast and continuous, it leads to prolonged sampling periods.
+These issues mean traditional single-agent reinforcement learning training approaches can be exceedingly time-consuming.
+Applying reinforcement learning in federated learning requires addressing these problems, as increasing participant numbers escalates agent training time.
+Therefore, optimizing the action space for reinforcement learning to expedite the agent training process is an essential challenge to address.
+Why opt for phased reinforcement learning?
+We take an example to explain this problem.
+Consider a robot learning to cook through reinforcement learning, with the process divided into washing, chopping, and cooking stages.
+The robot must master each stage to prepare a successful dish.
+Traditional reinforcement learning aims to identify the optimal action across all stages simultaneously; however, mastering the initial stage is essential before progressing.
+By adopting a phased learning approach, the robot sequentially masters each stage, streamlining the learning process and leading to more effective outcomes.
+Similarly, if malicious models are not initially filtered out, the agent’s trial-and-error costs in weight assignment for these models will increase.
+To resolve these issues, we propose a staged reinforcement learning algorithm (FedDRL).
+The contributions of the paper are as follows.
+  • We design a federated learning framework that employs reinforcement learning for model fusion, designed to select trustworthy clients and optimally assign model weights.
+  • We propose an adaptive client selection strategy based on the A2C algorithm, dynamically identifying and selecting trustworthy clients while excluding malicious ones from the model fusion process based on situational analysis.
+  • We propose an adaptive weight assignment method that adaptively adjusts the weights according to the quality of their uploaded models.
+  • We propose an adaptive weight assignment method that adaptively adjusts the model fusion weights according to the quality of their uploaded models.
   • We present five types of model fusion scenarios to validate the performance of each algorithm.
@@
 2.1 Federated Learning
-Research of federated learning focuses on solving three problems: how to improve generalization of the global model on the server side, how to personalize the model on the client side.Therefore, Federated learning can be divided into server-side and client-side optimization algorithms.
-Google first proposed the FedAvg method [1] to solve the server-side global model fusion problem .
-However, the client in the actual scenario has device and sample heterogeneity problems, which can lead to global model convergence problems in the FedAvg method.
-To optimize the global model convergence problem, Karimireddy et al.
-proposed the Scaffold algorithm [3], which accelerates the convergence of the global model by introducing a control variable to correct the client-side drift problem.
+Research in federated learning primarily aims to address two challenges: enhancing the generalization of the global model on the server side and personalizing the model on the client side.
+Consequently, federated learning algorithms are bifurcated into server-side and client-side optimization strategies.
+Google initially introduced the FedAvg algorithm [2] to address the problem of server-side global model fusion.
+To improve global model convergence, Karimireddy et al.
+developed the Scaffold method [3], which mitigates client-side drift by integrating a control variable.
+Similarly, Li et al.
+introduced FedProx [4], applying a regularization function to client models to correct deviations.
+Additionally, Wang et al.
+unveiled FedNova [5], addressing global model convergence issues by normalizing parameters on both client and server ends.
+Furthermore, Li et al.
+have introduced the MOON [6] technique, leveraging model comparison learning to enhance global model convergence.
+Chen et al.
+[29].
+also proposed a client identification method based on model parameter features to achieve trustworthy federated learning.
+While those approaches enhance the global model’s convergence speed, practical federated learning situations reveal variances in the quality of models trained by individual participants.
+These discrepancies stem from the diversity in computational resources and the calibre of data samples available to each participant.
+Additionally, variations arise due to the quantity and type of samples possessed by each participant, a phenomenon known as Non-IID (Non-Independent and Identically Distributed).
+Consequently, these factors complicate the attainment of optimal global model aggregation in the Non-IID environments.
+2.2 Challenges of Non-IID Data Distribution
+The Non-IID data issue significantly impacts federated learning models’ convergence.
+Zhao et al.
+explored various federated learning methods’ performance on non-IID datasets, demonstrating significant accuracy challenges [15].
+Accordingly, several studies have addressed the non-IID dilemma in federated learning.
+For instance, Zhang et al.
+proposed the FedPD approach [16], optimizing models and communication for non-convex objective functions.
+Moreover, Gong et al.
+introduced AutoCFL [17], utilizing a weighted voting client clustering strategy to mitigate non-IID and imbalanced data effects.
+Huang et al.
+developed FedAMP [18], which addresses Non-IID data-induced client-side model personalization issues through personalized model updates.
 Li et al.
-proposed the FedProx algorithm [2], which implements the correction of deviated clients by adding the regularization function to the client model.
-Wang et al.
-proposed the FedNova algorithm [4], which solves the global model convergence problem by normalizing the parameters on both the client and server sides.
-Li et al.
-also proposed the MOON [5] method based on model comparison learning, which solves the global model convergence problem by calculating the similarity between the global model and the client model and continuously decreasing the distance representation of both to achieve the training of the client-side model.
-2.2 The challenges of Non-IID Data Distribution
-The Non-IID problem is the key that leads to the convergence of the federated learning model.
-Zhao et al.
-study the performance of different federated learning methods on the Non-IID dataset and experimentally demonstrate that the Non-IID problem poses a significant challenge to the accuracy of the federated learning algorithm [13].
-Therefore some research focuses on solving the Non-IID problem of federated learning.
+devised Fedbn [19], incorporating a batch normalization layer into local models to address feature shift challenges due to data heterogeneity.
+Briggs et al.
+suggested a hierarchical clustering method (FL+HC) [1], improving Non-IID dataset model performance by grouping clients for independent model training.
+Additionally, Gao et al.
+offered the Feddc approach [20], bridging client and global model parameter disparities through a control variable.
+Lastly, Mu et al.
+introduced Fedproc [21], directing client model training by integrating a comparative loss between client and global models.
+Chen [7] et al.
+proposed a federated learning method based on adaptive knowledge distillation to improve the accuracy of heterogeneous model scenarios.
+Although these methodologies advance Non-IID issue mitigation in federated learning, they typically assign uniform fusion weights to all clients, failing to exclude malicious or low-quality model contributions.
+Consequently, dynamically selecting clients for fusion and adaptively calculating each model’s weight remains critical for successful global model integration.
+2.3 Federated Reinforcement Learning
+Given the adaptive learning potential of reinforcement learning, its application within federated learning contexts has garnered interest.
+Some research has concentrated on leveraging reinforcement learning to boost global model performance.
+For instance, Wang et al.
+introduced the Favor method [10], which adaptively selects clients for model fusion.
+Sun et al.
+developed the PG-FFL framework [22], addressing the challenge of client weight computation during model fusion.
+Additional studies have applied reinforcement learning for device optimization within federated IoT frameworks.
 For example, Zhang et al.
-proposed the FedPD method [14], which offers strategies that can handle non-convex objective functions to optimize models and communication.
-In addition, Gong et al.
-proposed the AutoCFL method [15], which adaptively adjusts the number of training rounds for the client while introducing weighted voting of the client clustering strategy to eliminate the effects of Non-IID and imbalanced data.
-Huang et al.
-proposed the FedAMP method [16], which solves the client-side model personalization problem caused by Non-IID data by implementing a personalized update of the client-side model.
-Li et al.
-proposed the Fedbn method [17], which adds a batch normalization layer to the local model to solve the problem of feature shift in data heterogeneity.
-Briggs et al.
-proposed a hierarchical clustering approach (FL+HC)[18], which optimizes the model’s performance on Non-IID datasets by training clients in groups and then having each group trained independently on a dedicated model.
-Furthermore, Gao et al.
-proposed the Feddc method [19], which addresses the data heterogeneity problem by bridging the gap between client and global model parameters through a control variable.
-Finally, mu et al.
-proposed the Fedproc method [20], which guides the client model training by introducing a loss of comparison between the client model and the global model, thus enabling the global model to perform well on Non-IID data to achieve good performance.
-2.3 Federated reinforcement learning
-Since reinforcement learning has the advantage of adaptive learning, some research applies reinforcement learning to federated learning.
-Some research focuses on improving the performance of global models based on reinforcement learning methods.
-Such as Wang et al.
-proposed the Favor method [7], which adaptively selects clients participating in the model fusion.
-Sun et al.
-also proposed the PG-FFL framework [6], which can solve the problem of calculating client weights during model fusion.
-Some work uses reinforcement learning to solve device optimization problems in a federated IoT environment.
-such as Zhang et al.
-uses the DDPG method [9] to achieve the selection of optimal devices.
-Zhang proposed the FedMarl algorithm [21], which uses Multi-agent reinforcement learning to achieve node selection in federated learning.
-Yang et al.
-also proposed a digital twin architecture (DTEI) [10], which uses reinforcement learning methods to solve the device selection problem.
-Some research focus on solving the resource optimization and scheduling problem in IoT environment.
-Zhang et al.
-proposed the RoF approach[11], which achieves optimal resource scheduling based on multi-intelligent reinforcement learning.
-Rjoub et al.
-designed a trusted device selection method [22] based on a reinforcement learning approach to achieve resource scheduling.
-Rjoub et al.
-proposed the DDQN-Trust method [12], which uses the Q-learning approach to evaluate the credit scores of devices to achieve optimal scheduling.
-To optimize the communication problem of federated learning, Yang et al.
-proposed a reinforcement learning-based model evaluation method [23] to participate in model training and model fusion by selecting the optimal devices.
-Most current research uses reinforcement learning to solve federated learning problems in IoT environments.
-These include optimal device selection, resource optimization, and communication optimization.
-Some work uses reinforcement learning methods to solve the client-side weight assignment problem in federated learning.
-However, most of these methods are based on single-agent training methods and need to address the training time overhead of reinforcement learning.
+utilized the DDPG algorithm [11] for optimal device selection.
+Zhang also formulated the FedMarl strategy [23], employing multi-agent reinforcement learning for node selection.
+Similarly, Yang et al.
+proposed a digital twin architecture (DTEI) [12], applying reinforcement learning for device selection issues.
+Other investigations have addressed resource optimization and scheduling challenges within IoT contexts, such as Zhang et al.’s RoF methodology [13], which leverages multi-intelligent reinforcement learning for optimal resource scheduling.
+Additionally, Rjoub et al.
+have developed trusted device selection techniques [24] and the DDQN-Trust method [14], utilizing Q-learning to assess devices’ credit scores for optimal scheduling.
+To ameliorate federated learning communication issues, Yang et al.
+introduced a reinforcement learning-based model evaluation method [25], selecting optimal devices for training and fusion.
+Nevertheless, while these efforts predominantly focus on IoT environment applications—such as device selection, resource optimization, and communication enhancement—they seldom address federated learning’s model weight calculation challenges.
+Therefore, Zhang et al.
+proposed the $R^{2}$ed framework [27], employing the DDPG reinforcement learning method for adaptive client weight calculation.
+Chen et al.
+[28] designed a task platform for implementing trustworthy federation learning.
+Although current research addresses the issue of weight allocation in federated learning, it often neglects the training efficiency of the agents.
+Therefore, optimizing the training efficiency of agents is a significant challenge that needs attention.
 3 Method
 3.1 Problem Definition
-In this section, we first analyze the problems of the current federated learning approach and then propose a solution to the problem.
-In the federation learning process, we aim to get a higher-accuracy global model by aggregating all clients’ local models through the server side.
-In the federated learning environment, we define n clients are involved in model fusion, and the client is denoted as $C_{i}$ where $C_{i}\in\left\{C_{1},C_{2},C_{3}\ldots C_{n}\right\}$.
+In this section, we scrutinize the prevailing challenges of the current federated learning approach and subsequently propose a solution.
+In federated learning, the objective is to get the global model by amalgamating local models from all clients through server-side aggregation.
+We define n clients as involved in model fusion, and the client is denoted as $C_{i}$ where $C_{i}\in\left\{C_{1},C_{2},C_{3}\ldots C_{n}\right\}$.
 Each client has a network model $M_{i}$, where $M_{i}\in\left\{M_{1},M_{2},M_{3}\ldots M_{n}\right\}$.
@@
 where $\theta_{i}\in\left\{\theta_{1},\theta_{2},\theta_{3}\ldots\theta_{n}\right\}$.
-Also, we define the server-side model aggregation process for each round as equation 1.
+Additionally, the server-side model aggregation process per round is defined as shown in equation 1:
 $\theta_{\text{global }}=\sum_{i=1}^{N}w_{i}\theta_{i},\text{ where }w_{i}=\frac{S_{i}}{\sum_{i=1}^{N}S_{i}},\quad w_{i}\geq 0,\sum_{i=1}^{N}w_{i}=1$ (1)
-Where $w_{i}$ is the weight of each client model parameter.
-The traditional method of calculating this weight in federal learning is determined by the ratio of the number of samples of the client to the total number of samples.
-This means that the larger the number of samples corresponding to the client, the larger the weight assigned.
-However, in the actual environment, due to the heterogeneity of the client data and the computing resources, it is impossible to determine the weights by the ratio of the number of samples for each client.
-Therefore, we need to solve two problems.
-Problem 1: The sample size of a client accounts for 20% of the total samples, but the accuracy of this client model is only 53%, which will reduce the accuracy of the global model if the traditional federated model fusion algorithm is used.
-Problem 2: If the client involved in the model fusion has malicious attacks and intentionally modifies the model output to make its model accuracy 10%, the global model accuracy will be seriously affected if the malicious models are selected for fusion using the traditional federated model fusion algorithm.
-To solve these two problems, we need an adaptive weight calculation method that can set the weight of the malicious model to 0, which means that malicious models are prohibited from participating in the fusion process.
-At the same time, we realize the adaptive assignment of weights to each client and try to increase the weights of high-quality client models so that the global model gets the optimal accuracy rate.
-If we use the single-agent reinforcement learning approach to solve the above problems, we face another issue again.
-On the one hand, the action space of the agent can be huge as the number of fused clients increases, which also leads to a long time required to train the agent model.
-On the other hand, a single agent method can only interact with a single environment, and that sampling process takes a long time.
-We solve these two problems in two steps by borrowing the idea of hierarchical reinforcement learning, which can solve the problem of the long training time of reinforcement learning.
-These two stages are selecting trusted clients and the optimal weight assignment process.
-Stage 1: In this phase, We need select K-trusted models from N models to participate in the global model fusion.
-However, we do not know which client uploaded a malicious model.
-To solve the problem, We use reinforcement learning to achieve dynamic selection and autonomous screening of client models as equation 2.
-$\left\{M_{a},M_{\mathrm{b}},\ldots M_{\mathrm{k}}\right\}\leftarrow\text{ SelectTrustedModel }\left(\left\{M_{1},M_{2},\ldots M_{\mathrm{n}}\right\}\right)$ (2)
-Stage 2: Based on step 1, we assign optimal weights to the trusted models to improve the accuracy of the global model and represent the process as equation 3.
-$\left\{W_{1},W_{2},\ldots W_{\mathrm{n}}\right\}\leftarrow\operatorname{AdaptCalculateWeight}\left(\left\{M_{a},M_{b},\ldots M_{\mathrm{k}}\right\}\right)$ (3)
-$AdaptCalculateWeight(.)$ is a weight-adaptive computation method. $W_{i}$ is the optimal computational weight for each client of the output.
-Why do we use a phased approach to reinforce learning?
-Let’s use an example to illustrate the reason for using it.
-We need a robot to learn the cooking process using a reinforcement learning approach.
-The cooking steps are divided into three stages: washing, chopping, and cooking.
-The robot needs to learn the best actions to cook a good dish.
-Traditional reinforcement learning is to find the best action that satisfies all three stages at the same time.
-However, if the robot does not wash the vegetables well, it cannot cook a good dish in the subsequent stages either.
-Therefore, it is a complete waste of time to try the next step of action if the optimal action of the previous stage is not learned.
-If the robot learns in multiple stages, it first learns the optimal action in the cleaning process and then proceeds to the next step of learning the optimal action.
-This will save time and make it easier to get the best solution.
-Therefore, if the malicious models are not screened and the weight assignment method is used for malicious models, it will increase the cost of trial and error time in the reinforcement learning process.
-Therefore, we screen the trusted clients for weight assignment by a staged approach.
-This can exclude the trial and error cost of malicious models involved in weight assignment, thus optimizing the action space.
-Also, we use distributed reinforcement learning methods to increase the sampling speed and thus reduce the time overhead of training reinforcement learning.
-We will describe the implementation of each step in detail in the next section.
-3.2 Trusted Client Selection Method
-We propose the FedDRL approach to solve the two problems in the previous section.
-In the first stage, we select the trusted clients to achieve the fusion of the models by excluding the malicious clients.
-In the second stage, we compute the optimal model weights to improve the accuracy of the global model.
-The flow of the FedDRL method is shown in Figure 1.
-We select trusted clients using the A2C algorithm.
-A2C is a synchronization algorithm achieved by improving the A3C algorithm [24].
-The selection process of a trusted client for each communication round of federated learning can be treated as a Markov decision process.
-The reinforcement learning modeling in this phase is described as follows.
-State: For the m-th communication round, we collect n models uploaded by each client and select k clients to participate in the fusion of global models.
-We define the $S^{t}$ as a select k client to fusion model in state t, which is shown as Equation 4, and the state includes the selected client models, the accuracy of each model, and the accuracy of the global model aggregated out by the selected client.
-$S^{t}=\left\{\theta_{1}^{t},\theta_{2}^{t},\theta_{3}^{t}\ldots\theta_{k}^{t},acc_{1}^{t},acc_{2}^{t},acc_{3}^{t}\ldots acc_{k}^{t},acc_{global}^{t}\right\}$ (4)
-Action: Our main objective is to eliminate the malicious models and select the trusted models to participate in the fusion.
-To achieve this purpose, we randomly choose several models to fusion the global model and evaluate the accuracy of global models.
-If a malicious model is selected to participate in the fusion, the accuracy of the global model will be very low.
-If only trusted clients are selected to participate in the fusion, the accuracy of the global model will also increase.
-We define the action of selecting a model as $a^{t}\in\left\{0,1\right\}$, where one means that the model is selected and 0 means that the model is not selected.
-Thus, n models correspond to the action space that can be expressed as equation 5.
-$A^{t}=\left\{a_{1}^{t},a_{2}^{t},a_{3}^{t}\ldots a_{n}^{t}\right\},\quad a^{t}\in\left\{0,1\right\}$ (5)
-Reward: We define the accuracy of selecting all models aggregated as $Acc_{all}$; we take this accuracy as the baseline and also represent the global model accuracy $Acc_{m}$ obtained by the m-th randomly selected client models, and calculate the reward value $Reward_{1}$ by the difference of $Acc_{m}-Acc_{all}$.
-Meanwhile, to avoid the intelligent body only, It can be realized that the agent selects all trusted clients as much as possible.
-The higher the number of models the agent sets, the higher the value of this reward.
-Therefore, the total reward is defined as equation 6.
-$Reward=\begin{cases}\alpha*(Acc_{m}-Acc_{all})+\beta*(Number),Acc_{m}>Acc_{all}\\ \alpha*(Acc_{m}-Acc_{all}),Acc_{m}<Acc_{all}\end{cases}$ (6)
-Where $\alpha$ and $\beta$ are fixed values that can be custom adjusted to balance the importance of Reward1 and Reward2, preventing one of the rewards from being too large and affecting agent training.
-The agent’s goal is to obtain the maximum long-term reward value based on the discount factor $\gamma$, and the process is expressed as equation 7.
-$R(\tau)=\sum_{\mathrm{t}=1}^{\mathrm{T}}\gamma^{t}\mathrm{r}_{\mathrm{t}}$ (7)
-Since we use an A2C distributed reinforcement learning-based approach, as shown in Figure 1, the A2C architecture consists of a central node and K workers.
-Each worker contains an Actor and a Critic network, where the Actor-network generates action A, and the Critic network evaluates the action and gives the corresponding reward.
+The $w_{i}$ is the fusion weight of each model parameter.
+Traditional Federated Learning typically employs a weighted average approach for computing model fusion weights, with each model’s weight determined by its corresponding client’s data sample size relative to the total.
+Thus, clients contributing more data exert a greater influence on the aggregated model.
+However, this method fails to consider the quality of each client’s model and the potential inclusion of malicious models in real-world scenarios.
+We illustrate the deficiencies of the traditional federated fusion algorithm through two scenarios:
+Scenario 1: A client’s data represents 20% of the total, yet its model’s accuracy is merely 53%.
+Employing the conventional federated fusion algorithm in this case would detrimentally impact the global model’s accuracy.
+Scenario 2: A client engaged in model fusion launches malicious attacks, intentionally skewing its model’s output to reflect a mere 10% accuracy.
+If such malicious models are incorporated through the standard fusion process, the accuracy of the global model would be severely compromised.
+Addressing these challenges necessitates an adaptive weight calculation strategy capable of nullifying malicious models by assigning them a weight of zero, thus excluding them from the fusion process.
+Concurrently, this approach should dynamically adjust the weights of each client’s model, prioritizing those of higher quality to enhance the global model’s overall accuracy.
+Adopting a single-agent reinforcement learning strategy to tackle these issues introduces new challenges.
+As the number of clients increases, so too does the agent’s action space, prolonging the training duration.
+Additionally, a single-agent framework is limited to interacting with just one environment, further extending the sampling period.
+We propose a bifurcated solution inspired by hierarchical reinforcement learning to mitigate these concerns, thereby streamlining the lengthy reinforcement learning training process.
+This solution comprises two primary stages: the selection of trustworthy clients and the assignment of optimal weights.
+Figure 1: The Process of FedDRL framework
+Stage 1: During this phase, the objective is to identify K trustworthy models from a pool of N for inclusion in the global model fusion.
+Identifying clients who have uploaded malicious models is challenging.
+We address this by employing reinforcement learning to dynamically select and autonomously screen client models, as delineated in equation 2.
+$\{M_{a},M_{\mathrm{b}},\ldots,M_{\mathrm{k}}\}\leftarrow\text{SelectTrustworthyModel}\left(\{M_{1},M_{2},\ldots,M_{\mathrm{n}}\}\right)$ (2)
+Stage 2: Building on the first step, we then allocate optimal weights to the verified models to bolster the global model’s accuracy, formalized in equation 3.
+$\{W_{1},W_{2},\ldots,W_{\mathrm{n}}\}\leftarrow\operatorname{AdaptCalculateWeight}\left(\{M_{a},M_{b},\ldots,M_{\mathrm{k}}\}\right)$ (3)
+Here, $AdaptCalculateWeight(.)$ signifies a method for adaptive weight computation, and $W_{i}$ represents the optimal computational weight assigned to each client’s output.
+3.2 A Trustworthy Federated Learning Approach Based on Staged Reinforcement Learning
+To address these challenges, we introduce a trusted federated learning framework anchored in staged reinforcement learning (FedDRL).
+This framework unfolds across two distinct phases.
+In the first phase, we propose an adaptive client selection strategy aimed at identifying and selecting trustworthy clients for participation in model fusion.
+Subsequently, in the second phase, we formulate a model weight assignment algorithm designed to dynamically allocate fusion weight values to models based on the prevailing fusion environment.
+The process is depicted in Figure 1.
+3.2.1 Adaptive client selection method
+Once we have defined the base elements of reinforcement learning, We use a distributed A2C approach to train the agent; A2C is an improved method-based A3C algorithm [26].
+Figure 1 shows the A2C architecture, which consists of a central node and $K$ workers.
+Each worker contains an Actor and a Critic network, where the actor network generates action, and the Critic network evaluates the action and gives the corresponding reward.
 Meanwhile, each worker independently interacts with the related environment to achieve sampling and training of the Actor and Critic networks.
 In addition, the Actor and Critic networks of the central node are used to synchronize the network information of each worker and to achieve the fusion and sharing of network parameters of multiple workers.
-Figure 1: The Algorithm flow of FedDRL
-Therefore our main objective is to train Actor and Critic networks.
+Therefore, our main objective is to train Actor and Critic networks.
 We define the Actor-network parameters as $\pi(\theta)$ and the Critic network parameters as $V(w)$.
-Our working process for each worker and the central node is as follows.
-The process is shown in algorithm 1.
-Step 1: First, each worker initializes the local network by pulling the global network model parameters.
-Then, both networks train their respective Actor and Critic networks by interacting with the environment independently.
-Finally, the trained networks are uploaded to the central node.
-Step 2: After the central node collects the network parameters uploaded by all workers, it updates the global model by weighted averaging.
-Finally, it sends it to each worker to update the network parameters.
+The process of the worker and the central node is as follows.
+Step 1: Each worker initializes the local network by pulling the global network model parameters from the centre node.
+Then, each worker trains the Actor and Critic networks by interacting with the environment independently.
+Finally, the two networks are uploaded to the central node.
+Step 2: After the central node collects the network parameters uploaded by all workers, it updates the global model by the weighted averaging method.
+Then, the server sends the two networks to each worker.
 Steps 1 and 2 are repeated according to the total number of times to obtain the final global model.
-The training process of the step 1 neutralization network is as follows. The gradient of the primary communication algorithm of the policy network is calculated as equation 8.
+The training process for the step 1 neutralization network is as follows: The gradient of the primary communication algorithm of the policy network is calculated as equation 4.
 $\nabla_{\theta}J(\theta)=\nabla_{\theta}\log\pi\left(a_{t}\mid s_{t};\theta\right)A\left(s_{t},a_{t};\mathrm{w}\right)$ (4)
@@
 $A\left(s_{t},a_{t};\theta_{v}\right)=\sum_{i=0}^{k-1}\gamma^{i}r_{t+i}+\gamma^{k}V\left(s_{t+k};\mathrm{w}\right)-V\left(s_{t};\mathrm{w}\right)$ (5)
-The Loss function of the Actor-network is calculated as in equation 10, and The Critic network is calculated as in equation 11.
+The Loss function of the actor network is calculated as in equation 6, and The Critic network is calculated as in equation 7.
 $\nabla_{\theta}J(\theta)=\nabla_{\theta}\log\pi\left(a_{t}\mid s_{t};\theta\right)\left(\sum_{i=0}^{k-1}\gamma^{i}r_{t+i}+\gamma^{k}V\left(s_{t+k};\mathrm{w}\right)-V\left(s_{t};\mathrm{w}\right)\right)$ (6)
@@
 $w\leftarrow w+\nabla_{w}J(w),\quad\theta\leftarrow\theta+\nabla_{\theta}J(\theta)$ (8)
-Finally, the network parameters of the central node are calculated by the weighted average method. The process is equation 13.
+Finally, each worker uploads the Actor and Critic network to the server. Then, the network parameters of the server are calculated using the weighted average method. The process is equation 9.
 $w_{global}=\frac{1}{n}{\textstyle\sum_{1}^{n}}w_{i},\theta_{global}=\frac{1}{n}{\textstyle\sum_{1}^{n}}\theta_{i},i\in[1,n]$ (9)
-Algorithm 1 FedDRL Stage 1 Algorithm Process
+When the parameters of the Actor and Critic networks in the central stage are updated, the central node sends down these two networks to all workers, and each worker uses the updated networks to continue interacting with the external environment.
+The process is repeated for the specified number of rounds until the agent at the central node can obtain a stable reward value.
+The process is shown in algorithm 1.
+Algorithm 1 The process of trustworthy client selection
 1: Client Models $\left\{m_{1}^{t},m_{2}^{t},m_{3}^{t},\ldots m_{\mathrm{n}}^{t}\right\}$, Round T, Worker Number K, Sampling Step Length S
@@
 24:Output Trusted Client Model List M = $\left\{m_{1}^{t},m_{2}^{t}\ldots m_{k}^{t}\right\}$
-3.3 Adaptive weights calculation method
-In this phase, we implement a distributed reinforcement learning approach based on TD3 [25].
+3.2.2 Adaptive model weight calculation method
 In this phase, our main objective is to achieve the optimal weight assignment for each model.
-Since the weight values of the client models correspond to a continuous action space, we use the TD3 algorithm to accomplish this process.
-For each communication round in federation learning, we assume that K plausible client models are screened in the previous stage, where the parameters of the client models are set.
-The number of samples corresponding to the clients is defined as.
-These two terms are fixed in each federal communication round.
-We need to assign the weight values corresponding to each model to the k models based on the quality of the models, which is defined.
-We validate the accuracy Acc of each model based on the retained data.
-We need to train the reinforcement learning intelligences in each communication round of federation learning and achieve the global model fusion by assigning the optimal model weights by this intelligences.
-The process is described as follows.
-(1) We define the t-th time, the action adopted by the agent to assign the weights of each model as equation 14.
+For each communication round, we assume that $K$ trustworthy client models were selected.
+We need to train the agents in each communication round and use the weight output of the agent to achieve the global model fusion.
+We first describe the process of global model fusion for agent-based actions.
+We define $\theta_{i}$ as the $i$-th client model, and the all client models as $\{\theta_{1},\theta_{2},\cdots,\theta_{k}\}$.
+We also define $s_{i}$ as the number of samples of $i$-th client.
+The process is as follows:
+(1) In this step, the agent needs to output the weight values for each model. We define the $t$-th time, the action adopted by the agent as equation 10.
 $\mathrm{W^{t}}=\left\{w_{1}^{t},w_{2}^{t},w_{3}^{t}\ldots w_{k}^{t}\right\}$ (10)
+$w_{i}$ is the $i$-th weight value output by agent for $i$-th model.
 (2) We aggregate the global models based on the model weights assigned by the agent, and the process is expressed as equation 11.
-$\theta_{global}^{k}=\sum_{\mathrm{i}=1}^{\mathrm{t}}w_{i}^{t}\theta_{i}$ (15)
-(3) We use the global model aggregated in step 2 to calculate the accuracy of this model and the reward value based on the accuracy rate.
-And the reward value is returned to the intelligence.
-The intelligence is trained by steps 1-3 until the reward value of the intelligence is stable.
-Finally, the optimal weights of each model are output by the intelligent body.
-We model the process as follows.
-State: We define the state of the teeth reinforcement learning as including the number of data samples S for each client, the accuracy Acc of each client’s model, and the global model Acc based on this aggregation.
-The process is as in 16.
+$\theta_{global}^{k}={\textstyle\sum_{t=1}^{T}}w_{i}^{t}\theta_{i}$ (11)
+We aim to train the agent so that it can output the optimal fusion weight values based on the quality of each model.
+To accomplish this goal, we first describe the basic elements of reinforcement learning as follows:
+Environment: The external environment is the server-side global model fusion module, which fuses the global model based on the actions output by the agent and then verifies the accuracy of the global model on the reserved dataset on the server side.
+Finally, the server side feeds back to the agent the corresponding reward and punishment values based on the accuracy of the global model.
+State: We define the agent’s state information to include the number of samples corresponding to each client, the accuracy of each client’s model, and the accuracy of the global model fused using the weights output by the agent.
+as shown in 12.
 $\mathrm{S^{t}}=\left\{s_{1}^{t},s_{2}^{t},s_{3}^{t}\ldots s_{k}^{t},acc_{1}^{t},acc_{2}^{t},acc_{3}^{t}\ldots acc_{k}^{t},acc_{global}^{t}\right\}$ (12)
-Action: In each stage, we need to reassign each model’s weights based on the model’s quality.
-The action space is shown as 17.
+The $acc_{global}^{t}$ is the accuracy of the global model fused using the weights output by the agent.
+Action: In each stage, the agent needs to assign each model’s weights based on the model’s quality.
+The action space is shown as 13.
 $a_{i}^{t}$ denotes the weight value assigned to the $i$-th client in the state t, while the sum of the corresponding weight values of all clients is 1.
 $\mathrm{A^{t}}=\left\{a_{1}^{t},a_{2}^{t},...a_{k}^{t}\right\},{\textstyle\sum_{1}^{k}}a_{i}^{t}=1,a_{i}^{t}\in(0,1)$ (13)
-Reward: We define the model accuracy aggregated using the weighted average method as $Acc_{all}$.
-Also, we define the global model accuracy obtained by the overall model weights assigned to the client for the mth time as $Acc_{m}$.
+Reward: We define the model accuracy aggregated using the average method as $Acc_{all}$, where each model weight is $\frac{1}{N}$.
+At the $m$-th time, we define the weight set output by the agent as $W$.
+Then, we use the weight set to fusion the global model, and we define the accuracy of the global model as $Acc_{m}$.
 We calculate the reward value by subtracting the difference of $Acc_{m}$ from $Acc_{all}$.
@@
 $\varphi$, $\phi$ denotes the reward and penalty factors, respectively.
-So the reward is defined as equation 18.
-$Reward=\begin{cases}\varphi*(Acc_{m}-Acc_{all}),Acc_{m}>Acc_{all}\\ \phi*(Acc_{m}-Acc_{all}),Acc_{m}<Acc_{all}\end{cases}$ (18)
-We use multiple parallel environments to accelerate the sampling to speed up the training agent in this phase.
-The stage is shown in figure 1.
+So, the reward is defined as equation 14.
+$Reward=\begin{cases}\varphi\cdot(Acc_{m}-Acc_{all}),Acc_{m}>Acc_{all}\\ \phi\cdot(Acc_{m}-Acc_{all}),Acc_{m}\leq Acc_{all}\end{cases}$ (14)
+When we have finished defining the basic elements, We implement a distributed reinforcement learning approach based on TD3 [8] to train the agent.
+The training process is shown in figure 1.
 This stage includes a central Learner and multiple Worker nodes.
@@
 Multiple workers interact with each environment independently, thus achieving parallel sampling to improve the sampling efficiency.
-After each worker collects a certain batch of samples, Learner trains the agent by taking a certain amount of sample data from the experience pool.
-We use the TD3 algorithm for this phase of the method design.
-The TD3 algorithm consists of six network models, including an Actor-network $P(w)$, two Critic networks $Q_{1}\left(\theta_{1}\right),Q_{2}\left(\theta_{2}\right)$, and a target Actor-network $P^{\prime}(w)$, two target Critic networks $Q^{\prime}_{1}\left(\theta_{1}\right),Q^{\prime}_{2}\left(\theta_{2}\right)$.
-The network results of the TD3 algorithm are shown in figure 1.
-The learner randomly draws N batches of sample data from the buffer pool every certain round for training the model.
+After each worker collects a certain batch of samples, the Learner trains the agent by taking a certain amount of sample data from the experience pool.
+The TD3 algorithm consists of six network models, including an Actor network $P(w)$, two Critic networks $Q_{1}\left(\theta_{1}\right)$,$Q_{2}\left(\theta_{2}\right)$, and a target Actor-network $P^{\prime}(w)$, two target Critic networks $Q^{\prime}_{1}\left(\theta_{1}\right),Q^{\prime}_{2}\left(\theta_{2}\right)$.
+Each network is shown in figure 1.
+The Learner randomly draws $N$ batches of sample data from the buffer pool every certain round to train the model.
 The training processes are as follows.
-The process is shown in algorithm 2
 […]
@@
 $y_{t}\leftarrow r+\gamma\operatorname{Min}(q1,q2)$ (15)
-(3) Select the action based on the Actor-network, input the state and action into the Critic network separately, and let these two networks output the corresponding prediction reward sum.
+(3) Select the action based on the actor network, input the state and action into the critical network separately, and let these two networks output the corresponding prediction reward sum.
 (4) Calculate the TD error. The calculation formula is as equation 16.
@@
 $\begin{array}[]{l}w^{\prime}\leftarrow\tau w+(1-\tau)w^{\prime}\\ \theta_{1}^{\prime}\leftarrow\tau\theta_{1}+(1-\tau)\theta_{1}^{\prime}\\ \theta_{2}^{\prime}\leftarrow\tau\theta_{2}+(1-\tau)\theta_{2}^{\prime}\end{array}$ (19)
-Algorithm 2 FedDRL Stage 2 Algorithm Process
+Repeating the above steps for the specified number of rounds, we will get the trained agent. Finally, we output the optimal value of each model through the agent. The process is shown in algorithm 2
+Algorithm 2 The process of model weight calculation
 1: Client Models $\left\{\theta_{1}^{t},\theta_{2}^{t},\theta_{3}^{t},\ldots\theta_{\mathrm{n}}^{t}\right\}$, Round R, Worker Number N, Buffer Memory Pool M Initialize Learner Parameters: Actor Parameter P$\left(w\right)$, Critic Network $Q_{1}\left(\theta_{1}\right),Q_{2}\left(\theta_{2}\right)$ Target Actor Parameter $P^{\prime}\left(w^{\prime}\right)$, Target Critic Network $Q_{1}^{\prime}\left(\theta_{1}^{\prime}\right),Q_{2}^{\prime}\left(\theta_{2}^{\prime}\right)$ $w^{\prime}\leftarrow w,\theta_{1}^{\prime}\leftarrow\theta_{1},\theta_{2}^{\prime}\leftarrow\theta_{2}$
@@
 17:     Every d Rounds:
-18:     Update Actor Network: $\nabla_{w}J\left(w\right)=N^{-1}\sum\nabla_{w}Q_{\theta_{1}}\left(s,a\right)\mid_{a=P\left({s}\right)}\nabla_{w}P\left({s}\right)$
+18:     Update Actor-Network: $\nabla_{w}J\left(w\right)=N^{-1}\sum\nabla_{w}Q_{\theta_{1}}\left(s,a\right)\mid_{a=P\left({s}\right)}\nabla_{w}P\left({s}\right)$
 19:     Update Target Critic Network: $\theta_{1}^{\prime}\leftarrow\tau\theta_{1}+\left(1-\tau\right)\theta_{1}^{\prime},\theta_{2}^{\prime}\leftarrow\tau\theta_{2}+\left(1-\tau\right)\theta_{2}^{\prime}$
-20:     Update Target Actor Network: $w^{\prime}\leftarrow\tau w+\left(1-\tau\right)w^{\prime}$
+20:     Update Target Actor-Network: $w^{\prime}\leftarrow\tau w+\left(1-\tau\right)w^{\prime}$
 21:After R Rounds, Save Trained Model
@@
 4 SYSTEM DESIGN
-We propose a trusted federated model fusion framework(FedDRL) to solve the client weight assignment problem.
-The main workflow is divided into five steps.
-Figure 2 shows the system architecture of FedDRL.
+To establish a reliable federated learning process, we developed a framework for trustworthy federated learning (FedDRL).
+The framework employs a staged reinforcement learning approach to achieve trustworthy federated learning.
+In the first stage, we train agents to accomplish the selection of trustworthy clients to participate in global model fusion.
+Then, in the second stage, we also use the trained agent to dynamically adjust the fusion weights of each model and finally realize the optimal global model fusion.
+The framework workflow consists of six steps, as shown in Figure 2.
+Step 1 (Local Model Training): Each client downloads the global model, initializes its parameters accordingly, and conducts model training using local private data.
+Step 2 (Upload Model): After local model training, each client uploads its model parameters to the server.
+Step 3 (Select Trustworthy Clients): Upon receiving client model parameters, the server employs the $SelectTrustClient(.)$ algorithm to train an agent.
+Subsequently, the trained agent selects trustworthy clients.
 Figure 2: The system architecture of FedDRL
-Step 1: First, each client downloads the global model, initializes the model parameters using the global model, and then trains the model using local private data.
-Step 2: After each client’s training, each client uploads its respective model parameters to the server side.
-Step 3: After receiving the client’s model parameters, the server side performs the trusted client selection and adaptive weight assignment in two stages according to the FedDRL algorithm.
-In stage 1, the server uses the SelectTrustClient(.) algorithm to train the agent, then updates each network parameter of the agent, and finally outputs the optimal action to achieve the selection of trusted clients.
-In stage 2, the server uses the AdaptCalculateWeight(.) algorithm to train the agent, then updates each network parameter of the agent, and finally achieves the optimal weight assignment for each client through the agent.
-Step 4: The server side fuses the client models based on the calculated optimal weights and finally fuses the global models.
-Step 5: The server side distributes the global model to each client.
-We set several communication rounds and repeat steps 1-5 until the global model converges. The process is shown in Algorithm 3.
+Step 4 (Assigning Model Weights): The server Utilizes models from trustworthy clients and performs global model fusion.
+It then employs the $AdaptCalculateWeight(.)$ algorithm to train an agent, which optimizes weight assignments for each client model.
+Step 5 (Fusing Global Model): The server fuses the global model using the calculated weights from the previous step.
+Step 6 (Distribute Global Model): The server disseminates the global model to all clients, initiating the subsequent federation task.
+The federation task is set to execute a specified number of communication rounds until the final global model is obtained. This process is shown in Algorithm 3.
 Algorithm 3 The FedDRL framework
@@
 4:for $C_{i}$ from 1 to $N$ do
-5:     $M_{i}\leftarrow\operatorname{GetGlobalModel}(round=i)$ // Get global model and init cleint model
+5:     $M_{i}\leftarrow\operatorname{GetGlobalModel}(round=i)$ // Get the global model and init client model
 6:     $M_{i}\leftarrow\operatorname{TrainLocalModel}(D_{i})$ // Train model $M_{i}$ based Dataset $\left\{D_{i}\right\}$
@@
 11:     Train the Stage 1 Agent
-12:     Update the $\operatorname{SelectTrustClient}\left(.\right)$ Algorithm parameters // According Algorithm 1
+12:     Update the $\operatorname{SelectTrustClient}\left(.\right)$ Algorithm parameters // According to Algorithm 1
 13:     $\left\{M_{a},M_{b},\ldots M_{\mathrm{k}}\right\}\leftarrow\operatorname{SelectTrustClient}\left(\left\{M_{1},M_{2},\ldots M_{n}\right\}\right)$
 14:     Train the Stage 2 Agent
-15:     Update the $\operatorname{AdaptCalculateWeight}\left(.\right)$ Algorithm parameters // According Algorithm 2
+15:     Update the $\operatorname{AdaptCalculateWeight}\left(.\right)$ Algorithm parameters // According to Algorithm 2
 16:     $\left\{W_{1},W_{2},\ldots W_{\mathrm{n}}\right\}\leftarrow\operatorname{AdaptCalculateWeight}\left(\left\{M_{a},M_{b},\ldots M_{k}\right\}\right)$
@@
 5.1.1 Experiment datasets
-We validate the FedDRL framework on three image classification datasets.
-Fashion-MNIST dataset: Fashion-MNIST consists of 70,000 samples in 10 classes.
-CIFAR-10 dataset: CIFAR-10 dataset consists of 60000 images in 10 classes.
-CIFAR-100 dataset: CIFAR-100 has 100 classes containing 600 images for each class.
-Data set partitioning: To simulate the distribution of non-IID datasets on the clients, we use the method in the paper [26] to partition different clients’ data on the open-source dataset using the Dirichlet function.
-The Dirichlet function could simulate the data partitioning of each client by setting the alpha parameter.
-When the alpha parameter tends to zero, the data distribution of clients is assigned to only a certain class of the public dataset; when the alpha parameter tends to positive infinity, the closer the distribution of client data, which is the amount and class of data is consistent across clients.
-To simulate the distribution of non-iid data, we set alpha to 1 to split the three datasets into ten clients.
-The data distribution is shown in Figure 3.
-Refer to caption (a) fashionmnist
+We evaluated the FedDRL framework using three distinct image classification datasets:
+Fashion-MNIST: This dataset includes 60,000 training samples and 10,000 test samples, each a 28x28 grayscale image, classified into one of 10 categories.
+CIFAR-10: The CIFAR-10 dataset comprises 60,000 32x32 colour images, evenly distributed across ten classes, each containing 6,000 images.
+CIFAR-100: Similar in size to CIFAR-10 but with a broader spectrum, CIFAR-100 features 100 classes with 600 images each, totalling 60,000 colour images.
+Data Set Partitioning: For simulating non-IID data distribution among clients.
+We utilized the Dirichlet function to segregate data across various clients in the open-source dataset.
+This method can partition the data for each client by adjusting the alpha parameter.
+As the alpha parameter approaches zero, clients’ data distributions are skewed towards specific classes within the dataset.
+Conversely, as alpha increases towards infinity.
+Using the CIFAR-10 dataset as a case study, we set alpha to 1, thereby dividing the three datasets among ten clients.
+In the figure, Different categories are represented by distinct colours, and the length of each segment within the graphs reflects the sample count within that category.
+The resulting data distribution is illustrated in Figure 3.
+Refer to caption (a) Fashion-MNIST
 Refer to caption (b) CIFAR-10
@@
 5.1.2 Comparison of Methods
-We compared the FedDRL algorithm with two traditional federated learning algorithms.
-FedAvg[1]: The FedAvg method is the baseline of federated learning, which allocates the weight of the client model by calculating the ratio of the number of samples from each client to the total number of samples.
-FedProx[2]: this algorithm improves the FedAvg method by adding a regularization term to the client model to optimize the performance of the federated learning.
+We contrasted the FedDRL algorithm with two established federated learning approaches.
+FedAvg[2]: Serving as the foundational benchmark in federated learning, the FedAvg method determines the weight of each client model based on the proportion of samples contributed by the client relative to the aggregate sample size.
+FedProx[4]: Enhancing the FedAvg approach, FedProx incorporates a regularization term within the client model, thereby refining federated learning performance.
 5.1.3 Experimental Metrics
-[…]
-$A_{\text{gobal }}=\left\{A_{\text{gobal }}^{(1)}\ldots A_{\text{gobal }}^{(t)}\ldots A_{\text{gobal }}^{(m)}\right\}$ (24)
-To demonstrate the behavior of each client. We also calculated the accuracy of the local model for each client in each round, which is expressed as 25.
-$A_{\text{c}}=\left\{A_{\text{c}}^{(1)}\ldots A_{\text{c}}^{(t)}\ldots A_{\text{c}}^{(m)}\right\},c\in\left[1,2...n\right]$ (25)
+We employed accuracy as the metric to gauge the performance of the global model in multi-classification tasks and across individual clients.
+Assuming n clients engage in model fusion with m communication rounds, the accuracy of the global model in the t-th round is denoted as $A_{\text{global}}^{(t)}$.
+The collective global model accuracies across all rounds are represented as follows:
+$A_{\text{global}}=\{A_{\text{global}}^{1},A_{\text{global}}^{2},\ldots,A_{\text{global}}^{m}\}$ (20)
+We denote the accuracy of the $c$-th client’s model as $A_{c}$. Additionally, we document each client model’s accuracy per round, compiling these as follows:
+$A_{\text{c}}=\{A_{\text{c}}^{1},A_{\text{c}}^{2},\ldots,A_{\text{c}}^{m}\},c\in[1,2,\ldots,n]$ (21)
 5.1.4 Experimental Configuration
-Hardware configuration: We completed the experiments using a high-performance workstation with an Intel i9-12900k processor, 64GB RAM, and an NVIDIA RTX3090 graphics card.
-Software configuration: We used two frameworks to complete the Federated Learning and Reinforcement Learning experiments.
-We use the FedBolt, our self-developed experimental framework, to conduct the Federated Learning experiments.
-We use this framework to simulate a Federated Learning experiment environment with a different number of clients and different data distributions.
-For the training process of the reinforcement learning model, we use the stablebaseline3 framework [27], based on which we design two algorithms for trusted client selection and model weight assignment.
-Federeted Learning Network setup: We designed different network models for different datasets.
-For the CIFAR-10 and CIFAR-100 datasets, we create a 6-layer CNN network to train the model.
-We designed a 4-layer MLP network for the FASHION-MNIST dataset.
-DRL Network setup: Since we use a staged reinforcement learning strategy, we use a distributed A2C algorithm in the first stage.
-Each worker and the central node include a 6-layer MLP for the Actor and a 6-layer MLP for the Critic Network.
-In the second stage, we used the TD3 algorithm and a 6-layer MLP network for each module of TD3.
+Hardware Configuration: The experiments were conducted on a workstation equipped with an Intel i9-12900k CPU, 64GB RAM, and an NVIDIA RTX3090 GPU.
+Software Configuration: We utilized two distinct frameworks for the Federated Learning and Reinforcement Learning experiments.
+Federated Learning trials were carried out using FedBolt, our custom-built framework, enabling simulation of varied client numbers and data distributions.
+For reinforcement learning model training, we employed the stablebaseline3 framework, designing two distinct algorithms for trusted client selection and model weight assignment.
+Network Setup: We implemented different network architectures tailored to each dataset.
+For CIFAR-10 and CIFAR-100, a 6-layer CNN was utilized for model training.
+Conversely, a 4-layer MLP was developed for the Fashion-MNIST dataset.
+Agent Network Setup: Implementing a staged reinforcement learning strategy necessitated the training of two distinct agents.
+The initial phase, adhering to Section 3.2.1, utilizes the A2C algorithm, with each worker and the central node comprising a 6-layer MLP actor and critic network.
+For the second phase, the TD3 algorithm outlined in Section 3.2.2 was employed for agent training, where each module within the TD3 setup incorporates a 6-layer MLP, with further details available in Section 3.2.2.
 5.2 Experimental Results
-We validate the FedDRL algorithm in three experiments.
-These include client attack experiments, low-quality model fusion experiments, and hybrid experiments.
-The client attack experiments focus on validating our proposed trusted client selection algorithm of stage 1.
-The low-quality model experiment mainly validates our proposed adaptive weight calculation algorithm of stage 2.
-The hybrid experiment contains both client attack and low-quality model components and mainly verifies the overall performance of the FedDRL method.
+We evaluate the FedDRL framework through four experiments: client attack scenarios, low-quality model fusion, hybrid scenarios, and multi-agent training efficiency.
+The client attack experiments assess the efficacy of the trustworthy client selection algorithm (stage 1).
+The low-quality model fusion experiment examines the adaptive weight calculation method (stage 2).
+The hybrid experiments, combining client attacks and low-quality model elements, validate the comprehensive performance of FedDRL.
+The final experiment focuses on the training efficiency of multi-agents.
 5.2.1 Malicious Client Attack Experiment
-We define three attack types to simulate the client-side attack behavior in federated learning.
-To accomplish the experiments, we simulated different numbers of clients and different types of attacks in three datasets.
-The experimental configuration of the client attack is shown in table 1.
+In this experiment, we define three types of client-side attacks in federated learning to evaluate our FedDRL framework under adversarial conditions.
+The experiment spans different client numbers and attack types across three datasets, detailed in Table 1.
 Type 1: The client directly uploads the initialized model or makes the model accuracy less than 10% by modifying the model’s hyperparameters.
 Type 2: We use falsified data to perform the attack.
-We use a certain percentage of forged data to participate in model training (e.g., mix the cifar10 dataset with $80\%$ of cifar100 data and generate these cifar100 data labels as cifar10 corresponding label types).
+We use a certain percentage of forged data to participate in model training (e.g., mix the CIFAR-10 dataset with $80\%$ of CIFAR-100 data and generate these CIFAR-100 data labels as CIFAR-10 corresponding label types).
 We conduct the attack by faking sample data to train the client’s local model, thus reducing the client model’s accuracy.
@@
 We also set that these clients alternately initiate the attack behavior.
-Table 1: Client Attack Experimental Settings for Different Numbers of Clients
+Table 1: Experimental setup for malicious attack scenarios
 Number Attack Type Malicious ID Number of samples Accuracy of models ($\leq$)
-Client=5 Type 1 Client1 7750 $A\leq 10\%$
+Clients=5 Type 1 Client1 7750 $A\leq 10\%$
 Type 2 Client1 7750 $10\%\leq A\leq 20\%$
 Type 3 Client1 7750 Attack round $A\leq 10\%$
-Client=10 Type 1 Client1,Client6 4222, 4938 $A\leq 10\%$
+Clients=10 Type 1 Client1,Client6 4222, 4938 $A\leq 10\%$
 Type 2 Client1,Client6 4222, 4938 $10\%\leq A\leq 20\%$
 Type 3 Client1,Client6 4222, 4938 Attack round $A\leq 10\%$
-Client=15 Type 1 Client1,Client6,Client11 3670, 3314, 4454 $A\leq 10\%$
+Clients=15 Type 1 Client1,Client6,Client11 3670, 3314, 4454 $A\leq 10\%$
 Type 2 Client1,Client6,Client11 3670, 3314, 4453 $10\%\leq A\leq 20\%$
 Type 3 Client1,Client6,Client11 3670, 3314, 4453 Attack round $A\leq 15\%$
-According to the experimental setup, we compared the FedDRL algorithm with the two baseline algorithms.
-In the attack experiments, we set the total number of communication rounds to 100 rounds, and each node performs local model training with 1 epoch.
-To show the attack behavior of each client and the accuracy of different algorithms in each attack scenario in more detail, we counted the accuracy of each client’s local model and the accuracy of the server-side global model in each communication round.
+According to the experimental setup, we compared the FedDRL algorithm with the FedAvg and FedProx.
+In the attack experiments, we set the total number of communication rounds to 100 rounds, and each client performs local model training with one epoch.
+To show the attack behavior of each client and the accuracy of different algorithms more detail, we counted the accuracy of each client’s local model and the accuracy of the server-side global model in each communication round.
 The specific experimental results are shown in table 2.
-Table 2: Global model accuracy for different number of client attack types
+Table 2: Accuracy of each algorithm under different malicious attack scenarios
 DataSet Method Clients=5 Clients=10 Clients=15
@@
 Ours 0.412 0.438 0.431 0.426 0.432 0.421 0.423 0.412 0.422
-To show the effect of the FedDRL algorithm on global model fusion at each communication round.
-We conducted experiments using the CIFAR10 dataset on 5, 10, and 15 clients.
-And We compared FedDRL with FedAvg and FedProx algorithms for the global model accuracy.
+To show the effect of the FedDRL algorithm on global model fusion at each communication round, we conducted experiments using the CIFAR10 dataset on 5, 10, and 15 clients.
+We compared FedDRL with the FedAvg and FedProx algorithms for global model accuracy.
 We analyze the experimental results for different numbers of client models and different client data.
-In attack type 1, our algorithm outperforms the FedAvg and FedProx algorithms on all three data sets.
-In malicious data attack type 2, our algorithm outperforms the FedAvg algorithm and slightly underperforms the FedProx algorithm alone.
+In attack type 1, In malicious data attack type 2, our algorithm outperforms the FedAvg algorithm and slightly underperforms the FedProx algorithm alone.
 In the attack type 3 scenario, our algorithm outperforms the comparison algorithm in most cases, especially when multiple malicious clients are involved in model fusion.
-To show the relationship between the global and client model’s accuracy in each attack scenario.
-We conducted more detailed experiments on the Cifar10 dataset.
-The experimental results are shown in Figure 4, Figure 5, and Figure 6.
-In attack type 1, our trained agent can dynamically select trusted clients for model fusion and eliminate malicious models from participating, so our algorithm has higher reliability.
-In attack type 2, our algorithm is lower than FedProx because the FedProx algorithm uses control parameters to force the models of all client nodes to converge to the global model, which will improve the accuracy of the global model by improving the accuracy of the malicious model to some extent.
-Our trained reinforcement learning agent will filter out low-accuracy models to participate in the fusion after several communication rounds.
-In attack type 3, the trained agent can dynamically adjust the trusted clients, so our algorithm works reliably even in the case of rotating attacks by malicious clients.
-Refer to caption (a) 5 clients type1 FedAvg
-Refer to caption (b) 5 clients type1 FedProx
-Refer to caption (c) 5 clients type1 FedDRL
-Refer to caption (d) 10 clients type1 FedAvg
-Refer to caption (e) 10 clients type1 FedProx
-Refer to caption (f) 10 clients type1 FedDRL
-Refer to caption (h) 15 clients type1 FedAvg
-Refer to caption (i) 15 clients type1 FedProx
-Refer to caption (j) 15 clients type1 FedDRL
-Figure 4: Global model accuracy in client attack type 1
-Refer to caption (a) 5 clients type3 FedAvg
-Refer to caption (b) 5 clients type3 FedProx
-Refer to caption (c) 5 clients type3 FedDRL
-Refer to caption (d) 10 clients type3 FedAvg
-Refer to caption (e) 10 clients type3 FedProx
-Refer to caption (f) 10 clients type3 FedDRL
-Refer to caption (h) 15 clients type3 FedAvg
-Refer to caption (i) 15 clients type3 FedProx
-Refer to caption (j) 15 clients type3 FedDRL
-Figure 5: Global model accuracy in client attack type 2
-Refer to caption (a) 5 clients type3 FedAvg
-Refer to caption (b) 5 clients type3 FedProx
-Refer to caption (c) 5 clients type3 FedDRL
-Refer to caption (d) 10 clients type3 FedAvg
-Refer to caption (e) 10 clients type3 FedProx
-Refer to caption (f) 10 clients type3 FedDRL
-Refer to caption (h) 15 clients type3 FedAvg
-Refer to caption (i) 15 clients type3 FedProx
-Refer to caption (j) 15 clients type3 FedDRL
-Figure 6: Global model accuracy in client attack type 3
+Refer to caption (a) 10 clients FedAvg
+Refer to caption (b) 10 clients FedProx
+Refer to caption (c) 10 clients FedDRL
+Refer to caption (d) 15 clients FedAvg
+Refer to caption (e) 15 clients FedProx
+Refer to caption (f) 15 clients FedDRL
+Figure 4: The accuracy of the global model for different number of client in attack type 1
+To show the relationship between the global and client model’s accuracy in each attack scenario. We conducted more detailed experiments on the Cifar10 dataset.
+In attack type 1, the global model accuracy plummets with increasing malicious clients under FedAvg and FedProx, dropping below 40% and 20% in 10 and 15 client setups, respectively.
+Conversely, FedDRL’s dynamic client selection maintains higher reliability.
+However, Our trained agent can dynamically select trusted clients for model fusion and eliminate malicious models from participating, so our algorithm has higher reliability.
+The experimental results are shown in Figure 4.
+In attack type 2, our algorithm is better than FedAvg but lower than FedProx.
+The FedProx algorithm uses control parameters to force the models of each client to converge to the global model, which will improve the global model’s accuracy by improving the malicious model’s accuracy to some extent.
+Our trained agent will filter out low-accuracy models to participate in the fusion after several communication rounds.
+The experimental results are shown in Figure 5.
+Refer to caption (a) 10 clients FedAvg
+Refer to caption (b) 10 clients FedProx
+Refer to caption (c) 10 clients FedDRL
+Refer to caption (d) 15 clients FedAvg
+Refer to caption (e) 15 clients FedProx
+Refer to caption (f) 15 clients FedDRL
+Figure 5: The accuracy of the global model for different number of client in attack type 2
+Refer to caption (a) 10 clients FedAvg
+Refer to caption (b) 10 clients FedProx
+Refer to caption (c) 10 clients FedDRL
+Refer to caption (d) 15 clients FedAvg
+Refer to caption (e) 15 clients FedProx
+Refer to caption (f) 15 clients FedDRL
+Figure 6: The accuracy of the global model for different number of client in attack type 3
+In attack type 3 scenarios, the FedAvg and FedProx algorithms experience significant fluctuations in global model accuracy due to alternating attack behaviors by malicious clients.
+Conversely, the agent within the FedDRL framework adaptively selects trusted clients, effectively excluding malicious entities from participating in model fusion, thereby enabling the FedDRL algorithm to operate with stability.
+The experimental results are shown in Figure 6.
 5.2.2 Low-quality Model Fusion Experiments
-To validate the performance of our proposed algorithm FedDRL, we performed validation on three datasets, Fashion-MNIST, CIFAR-10, and CIFAR-100.
-Since all three are open source, these data are high-quality data.
-Therefore each client used these sample data to train their respective models, and these models did not exhibit significant accuracy differences.
-We set up the following experiments to verify the impact of low-quality models involved in global model fusion in a real-world scenario.
-(1) To verify the impact of different numbers of customer nodes on the algorithm, we conducted experiments on three datasets with 5, 10, and 15 clients, respectively.
-We also set the parameter alpha=1 of Dirichlet to complete the dataset partitioning corresponding to each client.
-(2) We introduced a method for low-quality models to participate in the fusion.
-At the beginning of each round of federated learning, these clients initialize the local model using the global model.
-After several rounds of communication, when the accuracy of the local model reaches a certain accuracy, we simulate the participation of low-quality models in global model fusion by controlling the accuracy of these client models (e.g., model accuracy is kept between 50% and 60%).
-These clients upload low-quality models to the server side.
-Our experimental configuration of low-quality simulation for a different number of clients in table 3.
+In evaluating our FedDRL framework, we undertook validation using the Fashion-MNIST, CIFAR-10, and CIFAR-100 datasets.
+Given their open-source nature, these datasets are of high quality, leading to minimal variance in model accuracy among clients utilizing them directly.
+Thus, to simulate real-world conditions, we incorporated low-quality models into the global fusion process.
+We established a model accuracy threshold, ensuring that models uploaded by low-quality clients did not exceed this threshold in any communication round.
+Experiments were carried out on the three datasets, with client groups of varying sizes—5, 10, and 15—participating in the global model fusion.
+We applied a Dirichlet distribution with parameter alpha=1 to achieve dataset segmentation among clients.
+We set some clients to upload low-quality models; after several communication rounds, we controlled these client models’ accuracy in global fusion, ensuring it remained within the 40% to 55% range.
+Details of these low-quality model experiment configurations are specified in Table 3.
+The FedDRL algorithm was compared against the FedAvg and FedProx methods across 100 communication rounds, with each client executing one epoch of local model training.
+Results are summarized in Table 4.
 Table 3: Experimental settings for low-quality model experiments
 Number Dataset Low-quality Model ID Number of samples Accuracy of models ($\leq$)
-Client=5 Fashion-MINST Client1 9061 53%
+Clients=5 Fashion-MINST Client1 9061 53%
 CIFAR-10 Client1 7750 52%
 CIFAR-100 Client1 9278 22%
-Client=10 Fashion-MINST Client1,Client5 5071, 7245 51%, 52%
+Clients=10 Fashion-MINST Client1,Client5 5071, 7245 51%, 52%
 CIFAR-10 Client1,Client5 4222, 6039 50%, 54%
 CIFAR-100 Client1,Client5 4191, 5491 22%
-Client=15 Fashion-MINST Client1,Client5,Client10 4405, 3752, 1809 52%, 51%, 53%
+Clients=15 Fashion-MINST Client1,Client5,Client10 4405, 3752, 1809 52%, 51%, 53%
 CIFAR-10 Client1,Client5,Client10 3670, 3128, 1509 49%, 52%, 55%
 CIFAR-100 Client1,Client5,Client10 3073, 3494, 2910 22% 19% 23%
-We compared the FedDRL algorithm with both FedAvg and FedProx algorithms.
-We set the number of communication rounds to 100.
-Each client performs one epoch of local model training.
-The experimental results are shown in table 4.
-Table 4: Global model accuracy of Low-quality model experiments
+The FedDRL algorithm was compared against the FedAvg and FedProx methods across 100 communication rounds, with each client executing one epoch of local model training.
+Results are summarized in Table 4.
+Details of these low-quality model experiment configurations are specified in Table 3.
+Table 4: Accuracy of each algorithm for low-quality modeling experiments
 Method Fashion-MINST CIFAR-10 CIFAR-100
-Clients=5 Clients=10 Clients=15 Clients=5 Clients=10 Clients=15 Clients=5 Clients=10 Clients=15
+C=5 C=10 C=15 C=5 C=10 C=15 C=5 C=10 C=15
 FedAvg 0.857 0.858 0.841 0.705 0.664 0.602 0.386 0.373 0.365
@@
 Ours 0.885 0.887 0.884 0.725 0.706 0.698 0.422 0.418 0.407
-To demonstrate the FedAvg, FedProx, and FedDRL algorithms for global model fusion at each communication round.
-We have taken the CIFAR10 dataset as an example and conducted experimental comparisons with 5, 10, and 15 clients, respectively.
-The experimental results of each algorithm are shown in Figure 5.
-The comparison experiments of CIFAR-10 show that the performance of FedAvg and FedProx algorithms keeps decreasing as the number of low-quality models involved in fusion increases.
-Also, these two algorithms fluctuate up and down with the accuracy of the low-quality models due to the number of samples used by these two algorithms to calculate the global model.
-Therefore, when the weight value of the model is fixed, the change in the accuracy of the model affects the change in the accuracy of the global model.
-Meanwhile, FedDRL has a faster convergence speed than both FedAvg and FedProx algorithms because the FedDRL algorithm can adjust the corresponding weights adaptively according to the quality of each client model, which mitigates the impact of low-quality models on the accuracy of the global model and thus accelerates the convergence speed of the global model to a certain extent.
-Refer to caption (a) 5 clients FedAvg
-Refer to caption (b) 5 clients type2 FedProx
-Refer to caption (c) 5 clients FedDRL
-Refer to caption (d) 10 clients FedAvg
-Refer to caption (e) 10 clients FedProx
-Refer to caption (f) 10 clients FedDRL
-Refer to caption (g) 15 clients FedAvg
-Refer to caption (h) 15 clients FedProx
-Refer to caption (i) 15 clients FedDRL
-Figure 7: Global model accuracy in Low-quality model.
+Employing the CIFAR-10 dataset for illustrative purposes, we performed comparative analyses for setups with 10 and 15 clients, respectively; the findings are depicted in Figure 7.
+The experiments indicate that the accuracy of the FedAvg and FedProx methods deteriorates as the prevalence of low-quality models increases.
+This decline can be attributed to these algorithms’ reliance on sample count for determining the fusion weight values of the models, where the inclusion of low-quality models adversely impacts the global model’s accuracy.
+Conversely, FedDRL surpasses both methodologies in terms of global model convergence speed and accuracy.
+This is because FedDRL adaptively recalibrates the weights assigned to each client’s model based on quality, thereby diminishing the adverse effects of low-quality models on the global model’s accuracy and consequently hastening the global model’s convergence rate.
+Refer to caption (a) 10 clients FedAvg
+Refer to caption (b) 10 clients FedProx
+Refer to caption (c) 10 clients FedDRL
+Refer to caption (d) 15 clients FedAvg
+Refer to caption (e) 15 clients FedProx
+Refer to caption (f) 15 clients FedDRL
+Figure 7: The accuracy of a global model for different numbers of client in Low-quality scenario.
 5.2.3 Hybrid experiment
-In this section, we fuse two attack types (type 1 and type 3) and simulate a low-quality model to participate in the fusion.
-We verify The performance of the FedDRL algorithm in this hybrid Scenario.
-We also compare with FedAvg and FedProx methods.
-We used the cifar10 dataset and conducted experiments with 10 and 15 client cases.
-Client 1 only uploads the initialization model for each training, Client 6 simulates low-quality models to participate in fusion, client 10 or Client 11 initiates attack behavior in odd-numbered communication rounds, and even-numbered rounds participate in training normally.
-The remaining nodes normally participate in each round of Federated learning tasks.
-The detailed configuration of the experiments is shown in table 5.
-After 100 communication rounds, the global model accuracy of each algorithm is shown in table 6.
-Table 5: Hybrid Experimental Settings for Different Numbers of Clients
+In this section, we establish a hybrid scenario incorporating two types of attacking clients (type 1 and type 3) alongside clients submitting low-quality models.
+We assess the effectiveness of the FedDRL algorithm within this mixed scenario and benchmark it against the FedAvg and FedProx approaches.
+Employing the CIFAR-10 dataset, we set different numbers of clients (10,15) participating in global model fusion, respectively.
+Client 1 persistently uploads merely the initial model at each round.
+Client 6 emulates the submission of low-quality models for fusion, and Client 10 or 11 engages in attack behaviour during odd communication rounds but normally participates during even rounds.
+The remainder of the nodes contribute routinely to each cycle of the federated learning tasks.
+The experimental setup specifics are delineated in Table 5.
+Table 5: Experimental settings for hybrid scenarios
 Number Client ID Type Number of samples Model Accuracy
-Client=10 Client1 Attack Type 1 4222 $A\leq 10\%$
+Clients=10 Client1 Attack Type 1 4222 $A\leq 10\%$
 Client6 Low-quality Model 4938 $45\%\leq A\leq 50\%$
 Client10 Attack Type 3 3560 Attack round $A\leq 15\%$
-Client=15 Client1 Attack Type 1 3670 $A\leq 10\%$
+Clients=15 Client1 Attack Type 1 3670 $A\leq 10\%$
 Client6 Low-quality Model 3314 $45\%\leq A\leq 50\%$
 Client11 Attack Type 3 4453 Attack round $A\leq 15\%$
-Table 6: Global model Accuracy of Hybrid Experimental in different clients’ number
+Refer to caption (a) 10 clients FedAvg
+Refer to caption (b) 10 clients FedProx
+Refer to caption (c) 10 clients FedDRL
+Refer to caption (d) 15 clients FedAvg
+Refer to caption (e) 15 clients FedProx
+Refer to caption (f) 15 clients FedDRL
+Figure 8: Comparison of global model accuracy between different algorithms.
+After completing 100 communication rounds, we present the global model accuracy for each algorithm in Table 6.
+The comparative global model accuracies and individual client model accuracies per communication round, as determined by these three algorithms, are depicted in Figure 8.
+The experimental outcomes from the hybrid scenario reveal that the FedAvg and FedProx algorithms falter in properly conducting global model fusion due to the adversarial behavior of certain clients.
+Incorporating malicious models under traditional algorithmic frameworks significantly degrades the global model’s accuracy.
+Table 6: Accuracy of each algorithm for hybrid scenarios experiments
 Method Fashion-MINST CIFAR-10 CIFAR-100
@@
 Ours 0.876 0.883 0.701 0.698 0.426 0.418
-Refer to caption (a) 10 clients FedAvg
-Refer to caption (b) 10 clients FedProx
-Refer to caption (c) 15 clients FedDRL
-Refer to caption (d) 15 clients FedAvg
-Refer to caption (e) 15 clients FedProx
-Refer to caption (f) 15 clients FedDRL
-Figure 8: Comparison of global model accuracy between different algorithms.
-The hybrid scenario experiments show that the FedAvg and FedProx algorithms can no longer perform global model fusion properly due to the attack behavior of the client.
-Meanwhile, malicious models’ involvement in the global model fusion makes the accuracy of the global model based on the traditional algorithm very low.
-In the FedDRL algorithm, we adopt the two-stage method.
-In the first stage, we eliminate malicious models from participating in a fusion.
-We use an adaptive weight fusion strategy in the second stage, reducing low-quality models’ influence.
-Thus our algorithms work correctly even in this hybrid scenario.
-The global model and the model accuracy of each client for each communication round of these three algorithms are shown in Figure 8.
-5.2.4 Discussion
-We analyze the experimental results for different model fusion scenarios as well as for different numbers of client models.
-In the client-side attack experiments, under attack type 1, our algorithm works better than the FedAvg method on all three datasets.
-Meanwhile, as the number of malicious nodes increases, the global model accuracy of both FedAvg and FedProx algorithms shows a serious decrease.
-In contrast, our algorithm still maintains a smooth convergence trend.
-Under attack type 2, our algorithm is better than the FedAvg algorithm and slightly lower than the FedProx algorithm.
-Under attack type 3, our algorithm outperforms the edAvg and FedProx algorithms on all three datasets, which no longer work under rotating attacks with multiple malicious clients.
-In contrast, our algorithm can work reliably.
-In the low-quality model experiments, our algorithm outperforms the FedAvg and FedProx algorithms on all three datasets.
-In the hybrid experiments, the FedAvg and FedProx algorithms no longer work properly, and our algorithm can fuse the global models properly.
-Since the FedAvg and FedProx methods use the number of samples to calculate the model weights, they cannot adaptively adjust the responses of malicious and low-quality models.
-They thus cannot be applied to these attack scenarios.
-Our proposed method is based on staged reinforcement learning for malicious node removal and optimal weight assignment for trusted clients and thus can work reliably in various environments.
+The experimental outcomes show that FedAvg and FedProx’s global model accuracies suffer from malicious attacks due to their weighted average-based fusion, which doesn’t block harmful participants.
+Conversely, the FedDRL algorithm, through its two-stage approach, initially filters out malicious models from fusion and subsequently applies an adaptive weight strategy to diminish the impact of substandard models.
+Consequently, our algorithm maintains operational integrity even within this complex scenario.
+5.2.4 Agent Training Efficiency in the FedDRL Framework
+In this segment, our primary objective is to assess the training efficiency of agents within the FedDRL framework.
+To expedite the training process, we have implemented optimizations in two key areas.
+Initially, we adopted a distributed reinforcement learning methodology, enabling multi-agents to interact concurrently with the external environment.
+Concurrently, we introduced a memory cache module designed to prevent redundant sampling by multiple agents.
+Experimental Scenarios: Our investigation encompasses varied attack scenarios across two distinct datasets: Fashion-MNIST and Cifar-10.
+In each scenario, we involve a total of 10 and 15 clients in the federated task, including 2 and 3 malicious clients accordingly.
+Comparison Experiments: To ascertain the efficacy of the FedDRL framework, we initiated experiments featuring 1, 5, 10, and 20 agents.
+To guarantee the stability of the reward values acquired by the final agents, we designated the number of iterations for each experimental group to be 10,000, 15,000, 20,000, and 25,000, correspondingly.
+Experimental metrics: Our evaluation involves counting the iterations necessary for reinforcement learning to reach stable rewards across different agent counts.
+We employ a sliding window approach to compute the average reward, depicting the progression of rewards attained by the agents.
+We define $r_{t}$ the agent’s reward obtains in the $t$-th interaction and the sliding window as $W$.
+The formula for calculating the average reward is represented as equation 22:
+$\bar{R}=\frac{1}{W}{\textstyle\sum_{t=1}^{W}}r_{t}$ (22)
+Reward Parameter Setting: Our reward function comprises two components: the global model accuracy reward and the reward for the number of credible nodes.
+For this experiment, these parameters are set to 100 and 10, respectively.
+Table 7: The iterations of obtaining stable rewards for different numbers of agents
+Dataset Attack Type The number of agents
+N=1 N=5 N=10 N=20
+Fashion-MNIST Type-1 25000 20000 16000 8000
+Type-2 25000 21000 15000 9000
+CIFAR-10 Type-1 25000 20000 12000 10000
+Type-2 25000 20000 14000 9000
+Refer to caption (a) Fashion-MNIST (Attack Type 1)
+Refer to caption (b) Fashion-MNIST (Attack Type 2)
+Refer to caption (c) Cifar-10 (Attack Type 1)
+Refer to caption (d) Cifar-10 (Attack Type 2)
+Figure 9: Global model accuracy in 15 clients attack type 1
+Experimental Results: In accordance with our experimental setup, we recorded the reward values for each iteration of the agents, as detailed in Figure 9.
+We have systematically arranged this information into Table 7 for enhanced clarity regarding the actual iterations across different experiments.
+The data reveals a notable trend: The single agent does not get the optimal reward in some attack scenarios, because the single agent is easy to fall into the local optimal solution.
+Meanwile, an increase in agents correlates with reducing the iterations required to achieve a stable reward.
+However, this relationship is not strictly proportional because the multi-agent independently train their respective Actor and Critic networks.
+Each agent necessitates a distinct number of iterations to ensure the stability of its individual networks.
+Nevertheless, the simultaneous interaction of multiple agents with the environment markedly decreases the sampling time, demonstrating a clear trade-off between computational resources and time.
+This strategy underlines the significant computational resources required, highlighting a deliberate exchange of increased computational demand for reduced computational time.
 6 CONCLUSION
-To solve the problem of assigning client weights due to data heterogeneity in a federated learning environment.
-We propose FedDRL, an adaptive weight calculation method based on staged reinforcement learning.
-The method comprises two stages: selecting trusted clients and adaptive weight assignment.
-In the first stage, our approach can eliminate malicious client models from participating in the global model fusion.
-It also can dynamically select trusted client models to participate in the next stage of model weight assignment.
-In the second stage, Our method can adaptively calculate the corresponding weights based on the model quality of each client.
-Also, we propose a distributed reinforcement learning method to accelerate the agent’s training.
-Finally, we proposed five types of model fusion scenarios to validate our approach, and experiments show that our proposed algorithm can work reliably in all kinds of model fusion scenarios while maintaining the accuracy of the global model.
-Currently, our algorithm requires separate training of the agent according to the client fusion scenarios.
-In future work, we will extend the proposed algorithm to accommodate model fusion tasks for different task scenarios.
-We will also continue improving the algorithm’s performance to achieve more client tasks.
+To realize trustworthy federated learning, We propose a trusted reinforcement learning framework (FedDRL) based on staged reinforcement learning.
+The framework comprises two phases: selecting trusted clients and adaptive weight assignment.
+In the first phase, we design a reward strategy to train the agent, which allows the trained agent to exclude malicious client models from participating in model fusion based on the environment, and it also adaptively selects trustworthy clients for model fusion.
+In the second phase, we design a dynamic model weight calculation method, which can adaptively calculate the corresponding weights based on the model quality of each client.
+In addition, we propose a distributed reinforcement learning method to accelerate agent training.
+Finally, we design five model fusion scenarios to validate our approach, and the experiments show that our proposed algorithm can work reliably in various model fusion scenarios while maintaining global model accuracy.
+Although a multi-agent distributed reinforcement learning approach can accelerate the agent training process, it sacrifices computational resources for the computational time.
+In our future work, we will continue to explore more lightweight and trustworthy federated learning methods.
+We will also investigate more efficient reinforcement learning methods for credible federated learning.